### v1

# Subvarieties of pointed Abelian $\ell$-groups

Abstract
This paper provides a complete classification of the subvarieties and subquasivarieties of pointed Abelian lattice-ordered groups ($\ell$-groups) that are generated by chains. We present two complementary approaches to achieve this classification.
First, using purely $\ell$-group-theoretic methods, we analyze the structure of lexicographic products and radicals to identify all join-irreducible members of the lattice of subvarieties of positively pointed $\ell$-groups. We provide a novel equational basis for each of these subvarieties, leading to a complete description of the entire subvariety lattice. As a direct application, our $\ell$-group-theoretic classification yields an alternative, self-contained proof of Komori’s celebrated classification of subvarieties of MV-algebras.
Second, we explore the connection to MV-algebras via Mundici’s $\Gamma$ functor. We prove that this functor preserves universal classes, a result of independent model-theoretic interest. This allows us to lift the classification of universal classes of MV-chains, due to Gispert, to a complete classification of universal classes of totally ordered pointed Abelian $\ell$-groups. As a direct consequence, we obtain a full description of the corresponding lattice of subquasivarieties. These results offer a comprehensive structural understanding of one of the most fundamental classes of ordered algebraic structures.
Keywords: varieties, pointed Abelian $\ell$-groups, quasivarieties, universal classes, lexicographic product, axiomatization, Abelian logic, MV-algebras, Mundici functor, Komori classification
1 Introduction
Lattice-ordered groups ($\ell$-groups) represent a foundational class of algebraic structures with a rich history of celebrated results [18, 16, 15, 5, 11]. The field remains an active area of research, largely because the lattice of subvarieties of $\ell$-groups is notoriously complex and not yet fully understood [22]. Even the simplest case, the variety of Abelian $\ell$-groups, presents a deep mathematical challenge.
Another extremely important class of algebraic structures is that of MV-algebras. Originally introduced as the algebraic semantics for Łukasiewicz logic [29], these algebras have been the subject of extensive study for decades [7, 13, 23, 25]. One of the most essential tools in their study is the celebrated Mundici functor, a categorical equivalence between MV-algebras and Abelian $\ell$-groups with a strong unit. This functor provides a powerful bridge between the two fields, allowing many fundamental results in the theory of MV-algebras to be obtained by translating problems into the more established context of $\ell$-groups.
However, the theory of MV-algebras is in some respects richer than that of Abelian $\ell$-groups, primarily because its language contains an additional designated constant. A similar level of expressive power can be achieved in the theory of Abelian $\ell$-groups by augmenting their language with an additional constant, which gives rise to the class of pointed Abelian $\ell$-groups. As this article demonstrates, this class allows one to reason about MV-algebras while remaining entirely within the realm of Abelian $\ell$-groups. Furthermore, this class forms a variety, making it a natural and well-behaved extension of the class of Abelian $\ell$-groups with a strong unit, which is not even first-order definable.
In addition to its algebraic significance, our work is also strongly motivated by logic. Pointed Abelian $\ell$-groups serve as the algebraic models for pointed Abelian logic [8], a system that can be viewed as a meeting-point of Łukasiewicz logic [29] and Abelian logic [6, 24]. This paper provides a complete classification of all axiomatic and semilinear finitary extensions of pointed Abelian logic. Further details on the logical perspective can be found in [8].
The starting point for our investigation is Komori’s foundational classification of the varieties of MV-algebras [21]. This was later connected to the $\ell$-group setting by Young [28], who used the Mundici functor to establish a correspondence between these varieties and varieties of positively pointed Abelian $\ell$-groups. This paper generalizes and extends Young’s work in several key aspects:
  1. 1.
    First, we broaden the scope from positively pointed Abelian $\ell$-groups to the entire class of pointed Abelian $\ell$-groups.
  2. 2.
    Second, we develop our theory semantically from first principles within the theory of Abelian $\ell$-groups, rather than relying on the Mundici functor and the theory of MV-algebras. This contrasts with the more syntactic approach of [28] and allows us to prove more general theorems (e.g., compare our Theorem 3.10 with [28, Lemma 4.6]). As a final consequence of our framework, we derive the original Komori classification in Theorem 6.3.
  3. 3.
    Third, while the work in [28] established the correspondence between Abelian $\ell$-groups and MV-algebras, it did not provide axiomatizations for the corresponding equational classes. This paper fills that gap by providing such axiomatizations, analogous to Komori’s original description of the varieties of MV-algebras in [21].
  4. 4.
    Finally, we will use Gispert’s classification of universal classes of MV-chains and quasivarieties of MV-algebras generated by chains from [14] to provide the corresponding classifications for pointed Abelian $\ell$-groups.
The paper is structured as follows. Section 2 introduces the necessary preliminaries from universal algebra and the theory of Abelian $\ell$-groups. Section 3 develops the core technical results concerning lexicographic decompositions of finitely generated totally ordered $\ell$-groups, culminating in Theorem 3.10. Section 4 provides a semantic characterization of the varieties of Abelian $\ell$-groups, presenting novel and more elegant proofs for some known results (e.g., Lemma 4.8). The central contribution of the paper is presented in Section 5, which gives a complete classification and axiomatization of all subvarieties of pointed Abelian $\ell$-groups. Finally, Section 6 applies our results to the theory of MV-algebras, providing an alternative proof of Komori’s classification. Moreover, by using Gispert’s classification of universal classes of MV-chains, we provide a similar classification for $\ell$-group chains and consequently obtain a classification of all quasivarieties generated by $\ell$-group chains.
2 Preliminaries
In this section we introduce some basic terminology and well-known results about Abelian $\ell$-groups. We assume that the reader has a basic background in universal algebra. No advanced background is assumed throughout the paper. We will use the standard notations $\mathbf{H},\mathbf{I},\mathbf{S},\mathbf{P},\mathbf{P}_{\mathrm{U}}$ to denote closure under homomorphisms, isomorphisms, subalgebras, products, and ultrapowers, respectively. We will write ${\boldsymbol{\mathit{A}}}\in{{\mathbb{K}}}_{SI}$ iff the algebra $\boldsymbol{\mathit{A}}$} is a subdirectly irreducible element of the class $\mathbb{K}$}}.
First let us recall Birkhoff’s Theorem and its variants.
Theorem 2.1 (Birkhoff [3, 4]).
Let $\mathbb{K}$}} be a class of algebras of the same signature. The following conditions are equivalent:
  1. 1.
    ${{\mathbb{K}}}=\mathbf{HSP}({{\mathbb{K}}})$.
  2. 2.
    ${{\mathbb{K}}}=\mathbf{ISP}({{\mathbb{K}}}_{SI})$.
  3. 3.
    ${{\mathbb{K}}}=\mathbf{ISPP}_{\mathrm{U}}(\{{\boldsymbol{\mathit{A}}}\in{{\mathbb{K}}}_{SI}\mid{\boldsymbol{\mathit{A}}}\text{ is finitely generated}\})$.
  4. 4.
    $\mathbb{K}$}} is the class of all models of some theory, all of whose axioms are equations.
Moreover, if algebras in $\mathbb{K}$}} have a group reduct, we can add the following:
  1. 5.
    $\mathbb{K}$}} is the class of all models of some theory, all of whose axioms are equations with the right side equal to 0.
Let $\boldsymbol{\mathit{A}}$} be an algebra, $\mathrm{(eq)}$ an equation, and $e$ an evaluation. We write ${\boldsymbol{\mathit{A}}}\vDash_{e}\mathrm{(eq)}$ to denote that $\mathrm{(eq)}$ is satisfied in $\boldsymbol{\mathit{A}}$} under the evaluation $e$. Conversely, ${\boldsymbol{\mathit{A}}}\nvDash_{e}\mathrm{(eq)}$ indicates that $\mathrm{(eq)}$ is not satisfied under $e$.
An equation $\mathrm{(eq)}$ is said to hold in an algebra $\boldsymbol{\mathit{A}}$}, denoted ${\boldsymbol{\mathit{A}}}\vDash\mathrm{(eq)}$, if it is satisfied for all possible evaluations. Conversely, if there is at least one evaluation for which $\mathrm{(eq)}$ is not satisfied, it does not hold in $\boldsymbol{\mathit{A}}$}, and we write ${\boldsymbol{\mathit{A}}}\nvDash\mathrm{(eq)}$.
We now provide the definition of an essential notion used in this paper.
Definition 2.2.
Let ${{\mathbb{K}}}\cup\{{\boldsymbol{\mathit{A}}}\}$ be a class of algebras of the same signature. Let $F$ be a finite subset of $A$ and ${\boldsymbol{\mathit{B}}}\in{{\mathbb{K}}}$. We say that a mapping $f_{F}:F\rightarrow B$ is a partial embedding of the set $F$ from $\boldsymbol{\mathit{A}}$} into $\boldsymbol{\mathit{B}}$} if $f_{F}$ is a one-to-one mapping such that for each $a_{1},\dots,a_{n}\in F$ and for each $\lambda\in\mathcal{L}$ such that $\lambda^{{\boldsymbol{\mathit{A}}}}(a_{1},\dots,a_{n})\in F$ we have
$f_{F}(\lambda^{{\boldsymbol{\mathit{A}}}}(a_{1},\dots,a_{n}))=\lambda^{{\boldsymbol{\mathit{B}}}}(f_{F}(a_{1}),\dots,f_{F}(a_{n})).$
We say that $\boldsymbol{\mathit{A}}$} is partially embeddable into the class $\mathbb{K}$}} if for each finite set $F\subseteq A$ there is a ${\boldsymbol{\mathit{B}}}\in{{\mathbb{K}}}$ and there is a partial embedding $f_{F}$ of $F$ into $\boldsymbol{\mathit{B}}$}. If $\boldsymbol{\mathit{A}}$} is partially embeddable into $\{{\boldsymbol{\mathit{B}}}\}$ we just say $\boldsymbol{\mathit{A}}$} is partially embeddable into $\boldsymbol{\mathit{B}}$}.
Lemma 2.3.
Let $\mathbb{K}$}} be a class of algebras of the same signature. The following conditions are equivalent:
  1. 1.
    $\mathbb{K}$}} is the class of all models of some theory, all of whose axioms are universal formulas.
  2. 2.
    ${{\mathbb{K}}}=\mathbf{ISP}_{\mathrm{U}}({{\mathbb{K}}})$.
  3. 3.
    $\mathbb{K}$}} is closed under partial embeddings (i.e. for any algebra $\boldsymbol{\mathit{A}}$} that is partially embeddable into $\mathbb{K}$}} we get ${\boldsymbol{\mathit{A}}}\in{{\mathbb{K}}}$).
Now we will focus on Abelian $\ell$-groups.
Definition 2.4.
An algebra ${\boldsymbol{\mathit{A}}}=\langle{A,+,-,\lor,\land,0}\rangle$ is an Abelian $\ell$-group if $\langle{A,+,-,0}\rangle$ is an Abelian group, $\langle{A,\lor,\land}\rangle$ is a lattice and $\boldsymbol{\mathit{A}}$} satisfies the monotonicity condition, that is, for each $x,y,z\in A$ we get $x\leq y$ implies $x+z\leq y+z$.
It is well-known (see, e.g. [12, Chapter V]), that the defining conditions of Abelian $\ell$-groups can be expressed by means of equations, so they form a variety which we denote by $\mathbb{AL}$. Also, it is well-known that all Abelian $\ell$-groups are torsion-free.
We will denote by $\boldsymbol{\mathit{Z}}$} and $\boldsymbol{\mathit{R}}$} the $\ell$-groups of integers and reals with the underlying universes $\mathbb{Z}$ and $\mathbb{R}$.
The following results are also well-known.
Theorem 2.5 (Clifford [9]).
An Abelian $\ell$-group is subdirectly irreducible iff it is totally ordered.
Theorem 2.6 (Gurevich, Kokorin [17]).
All totally ordered Abelian $\ell$-groups are universally equivalent. Equivalently, for every totally ordered Abelian $\ell$-group $\boldsymbol{\mathit{A}}$} we have $\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{A}}})=\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}})$.
From Theorems 2.5 and 2.6 one can easily derive the following:
Theorem 2.7 (Khisamiev [20]).
The quasivariety of all Abelian $\ell$-groups is generated by $\boldsymbol{\mathit{Z}}$}. Equivalently, $\mathbf{ISPP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}})=\mathbb{AL}$.
Theorem 2.8 (Hölder’s Theorem [19]).
The following are equivalent for any Abelian $\ell$-group $\boldsymbol{\mathit{A}}$}:
  1. 1.
    $\boldsymbol{\mathit{A}}$} embeds into $\boldsymbol{\mathit{R}}$}.
  2. 2.
    $\boldsymbol{\mathit{A}}$} is Archimedean.
  3. 3.
    $\boldsymbol{\mathit{A}}$} is simple.
Lemma 2.9.
Let $\boldsymbol{\mathit{A}}$} be an $\ell$-subgroup of $\boldsymbol{\mathit{R}}$}. Then $A$ is a dense subset of $\mathbb{R}$ (with respect to the standard topology on $\mathbb{R}$) or ${\boldsymbol{\mathit{A}}}\cong{\boldsymbol{\mathit{Z}}}$.
Proof.
Since multiplication by any element of $\mathbb{R}$ is an $\ell$-group automorphism on $\boldsymbol{\mathit{R}}$} we can without loss of generality assume $1\in{\boldsymbol{\mathit{A}}}$. If $\boldsymbol{\mathit{A}}$} contains ${\boldsymbol{\mathit{Z}}}[\xi]$ for some $\xi\in\mathbb{R}\setminus\mathbb{Q}$, it is easy to show (using Euclidean algorithm on $1$ and $\xi$) that $A$ must be a dense subset of $\mathbb{R}$.
Otherwise, ${\boldsymbol{\mathit{A}}}\in\mathbf{S}({\boldsymbol{\mathit{Q}}})$. By [1, Theorem 2] and [1, Corollary 2] for every non-cyclic ${\boldsymbol{\mathit{A}}}\in\mathbf{S}({\boldsymbol{\mathit{Q}}})$ for each $\epsilon\in\mathbb{R}$ there is $a\in A$ such that $0<a\leq\epsilon$. Since $\boldsymbol{\mathit{A}}$} is also closed under addition, AA has to be a dense subset of $\mathbb{R}$. If $\boldsymbol{\mathit{A}}$} is cyclic then ${\boldsymbol{\mathit{A}}}\cong{\boldsymbol{\mathit{Z}}}$. ∎
Definition 2.10.
For an Abelian $\ell$-group ${\boldsymbol{\mathit{A}}}=\langle{A,+,-,\lor,\land,0}\rangle$ and $a\in A$ we define a pointed Abelian $\ell$-group ${\boldsymbol{\mathit{A}}}_{a}=\langle{A,+,-,\lor,\land,0,\mathtt{f}}\rangle$, [^3] It should be stressed that the choice of the symbol $\mathtt{f}$ has no algebraic motivation; its origin is purely logical. The symbol is chosen to align with its use as a ’falsum’ constant (representing falsehood) in logical systems. This becomes relevant when we consider pointed Abelian $\ell$-groups as algebraic counterparts to expansions of Abelian logic, as detailed in [24, 8]. where $\mathtt{f}^{{\boldsymbol{\mathit{A}}}_{a}}=a$. We denote the class of pointed Abelian $\ell$-groups by $p\mathbb{AL}$. We say that ${\boldsymbol{\mathit{A}}}_{a}\in p\mathbb{AL}$ is positively pointed if $a\geq 0$, negatively pointed if $a\leq 0$, and 0-pointed if $a=0$. We denote these classes by $p\mathbb{AL}^{+},p\mathbb{AL}^{-}$ and $p\mathbb{AL}^{0}$.
It is well known that congruences on Abelian $\ell$-groups coincide with convex subgroups, see e.g. [12].
Despite the fact that the multiplication symbol $\cdot$ is not present in the language of Abelian $\ell$-groups we will commonly use it in the traditional meaning of the iterated addition. Clearly, the classes $p\mathbb{AL},p\mathbb{AL}^{+},p\mathbb{AL}^{-}$ and $p\mathbb{AL}^{0}$ are varieties. Let us note that for an Abelian $\ell$-group $\boldsymbol{\mathit{A}}$} and $a\in A$ it holds that $\boldsymbol{\mathit{A}}$} is subdirectly irreducible iff ${\boldsymbol{\mathit{A}}}_{a}$ is subdirectly irreducible. Therefore, using Theorem 2.5 we obtain that a pointed Abelian $\ell$-group is subdirectly irreducible iff it is totally ordered. Another important fact is that the variety $p\mathbb{AL}^{0}$ is obviously term equivalent to $\mathbb{AL}$. The following can be proved.
Lemma 2.11.
  1. 1.
    $p\mathbb{AL}_{SI}=\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{-1},{\boldsymbol{\mathit{R}}}_{0},{\boldsymbol{\mathit{R}}}_{1})$.
  2. 2.
    $p\mathbb{AL}_{SI}^{+}=\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{0},{\boldsymbol{\mathit{R}}}_{1})$.
  3. 3.
    $p\mathbb{AL}_{SI}^{-}=\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{-1},{\boldsymbol{\mathit{R}}}_{0})$.
  4. 4.
    $p\mathbb{AL}_{SI}^{0}=\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{0})=\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{0})$.
  5. 5.
    $p\mathbb{AL}=\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{-1},{\boldsymbol{\mathit{R}}}_{0},{\boldsymbol{\mathit{R}}}_{1})$.
  6. 6.
    $p\mathbb{AL}^{+}=\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{0},{\boldsymbol{\mathit{R}}}_{1})$.
  7. 7.
    $p\mathbb{AL}^{-}=\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{-1},{\boldsymbol{\mathit{R}}}_{0})$.
  8. 8.
    $p\mathbb{AL}^{0}=\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{0})=\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{0})$.
Proof.
Let us recall that pointed Abelian $\ell$-groups are subdirectly irreducible iff they are linearly ordered.
All the inclusions $\supseteq$ are trivial. We prove the remaining inclusions in a convenient order.
  1. 4.
    Follows, since the class $p\mathbb{AL}_{SI}^{0}$ is term equivalent to $\mathbb{AL}_{SI}.$
  2. 1.
    Let us fix an arbitrary ${\boldsymbol{\mathit{A}}}_{a}\in p\mathbb{AL}_{SI}$. Using Theorem 2.6 and [8, Lemma 4.3] we obtain ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{ISP}_{\mathrm{U}}(\{{\boldsymbol{\mathit{R}}}_{b}\mid b\in\mathbb{R}\})$. Since for $b\in\mathbb{R}$ we have ${\boldsymbol{\mathit{R}}}_{b}\cong{\boldsymbol{\mathit{R}}}_{-1}$, ${\boldsymbol{\mathit{R}}}_{b}\cong{\boldsymbol{\mathit{R}}}_{0}$, or ${\boldsymbol{\mathit{R}}}_{b}\cong{\boldsymbol{\mathit{R}}}_{1}$ we obtain ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{-1},{\boldsymbol{\mathit{R}}}_{0},{\boldsymbol{\mathit{R}}}_{1})$.
  3. 2.
    For ${\boldsymbol{\mathit{A}}}_{a}\in p\mathbb{AL}_{SI}^{0}$ we already know ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{0},{\boldsymbol{\mathit{R}}}_{1})$. Therefore, we fix arbitrary ${\boldsymbol{\mathit{A}}}_{a}\in p\mathbb{AL}_{SI}^{+}\setminus p\mathbb{AL}_{SI}^{0}$. By the previous point we know that ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{-1},{\boldsymbol{\mathit{R}}}_{0},{\boldsymbol{\mathit{R}}}_{1})$. By [2, Theorem 5.6] we have
    $\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{-1},{\boldsymbol{\mathit{R}}}_{0},{\boldsymbol{\mathit{R}}}_{1})=\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{-1})\cup\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{0})\cup\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{1}).$
    Since ${\boldsymbol{\mathit{A}}}_{a}\in p\mathbb{AL}_{SI}^{+}\setminus p\mathbb{AL}_{SI}^{0}$, it follows $a>0$. Therefore, ${\boldsymbol{\mathit{A}}}_{a}\notin\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{-1})$ since ${\boldsymbol{\mathit{A}}}_{a}\nvDash\mathtt{f}\leq 0$. Thus ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{0},{\boldsymbol{\mathit{R}}}_{1})$.
  4. 3.
    Can be proved similarly as the previous point.
The other four points follow directly by applying Theorem 2.1. ∎
We will later prove a strengthening of the second part of Lemma 2.11 in Lemma 3.7.
3 The lexicographic decompositions
This section focuses on understanding the structure of the lexicographic product of $\ell$-groups. The lexicographic product is a key operation on $\ell$-groups. It allows for the creation of otherwise unintuitive $\ell$-groups that play an essential role in various classifications (see e.g. the famous Hahn Theorem from [18]). Since we will need to use this tool frequently in the following chapters, here we establish several basic properties of this construction. The main result of this section is Theorem 3.10, which is a stronger version of [28, Lemma 4.6], which tells us that from the viewpoint of universal classes of totally ordered Abelian $\ell$-groups we are not interested in $\ell$-groups without a strong unit. We will proceed to the definition of the lexicographic product.
Definition 3.1.
Let $\boldsymbol{\mathit{A}}$} be a totally ordered Abelian $\ell$-group and $\boldsymbol{\mathit{B}}$} be an Abelian $\ell$-group. We define the Abelian $\ell$-group ${\boldsymbol{\mathit{A}}}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}$ as the Abelian $\ell$-group ${\boldsymbol{\mathit{A}}}\times{\boldsymbol{\mathit{B}}}$ with modified lattice operations:
$\langle{a_{1},b_{1}}\rangle\lor\langle{a_{2},b_{2}}\rangle=\begin{cases}\langle{a_{1},b_{1}}\rangle&a_{1}>a_{2}\\ \langle{a_{2},b_{2}}\rangle&a_{1}<a_{2}\\ \langle{a_{1},b_{1}\lor b_{2}}\rangle&a_{1}=a_{2}\end{cases}$
and
$\langle{a_{1},b_{1}}\rangle\land\langle{a_{2},b_{2}}\rangle=\begin{cases}\langle{a_{1},b_{1}}\rangle&a_{1}<a_{2}\\ \langle{a_{2},b_{2}}\rangle&a_{1}>a_{2}\\ \langle{a_{1},b_{1}\land b_{2}}\rangle&a_{1}=a_{2}\end{cases}$
Note that for totally ordered Abelian $\ell$-groups ${\boldsymbol{\mathit{A}}},{\boldsymbol{\mathit{B}}}$ and Abelian $\ell$-group $\boldsymbol{\mathit{C}}$} it holds that ${\boldsymbol{\mathit{A}}}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}$ is totally ordered and moreover $({\boldsymbol{\mathit{A}}}\overrightarrow{\times}{\boldsymbol{\mathit{B}}})\overrightarrow{\times}{\boldsymbol{\mathit{C}}}\cong{\boldsymbol{\mathit{A}}}\overrightarrow{\times}({\boldsymbol{\mathit{B}}}\overrightarrow{\times}{\boldsymbol{\mathit{C}}})$. Therefore, we will commonly omit parentheses and we will just write ${\boldsymbol{\mathit{A}}}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}\overrightarrow{\times}{\boldsymbol{\mathit{C}}}$.
Lemma 3.2.
Let $\boldsymbol{\mathit{A}}$} be a finitely generated totally ordered Abelian $\ell$-group and $\boldsymbol{\mathit{B}}$} be a convex $\ell$-subgroup of $\boldsymbol{\mathit{A}}$}. Then ${\boldsymbol{\mathit{A}}}\cong({\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}})\overrightarrow{\times}{\boldsymbol{\mathit{B}}}$.
Proof.
Since $\boldsymbol{\mathit{A}}$} is finitely generated, also ${\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}}$ is finitely generated. Since ${\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}}$ is a finitely generated torsion-free Abelian group, it is group-isomorphic (but not necessarily $\ell$-group isomorphic) to a free Abelian group (see [27, Theorem 10.19]). Thus the group exact sequence $0\rightarrow{\boldsymbol{\mathit{B}}}\rightarrow^{\iota}{\boldsymbol{\mathit{A}}}\rightarrow^{\pi}{\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}}\rightarrow 0$ splits and there is a group homomorphism $p:{\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}}\rightarrow{\boldsymbol{\mathit{A}}}$ such that $\pi\circ p=id_{{\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}}}$. It is well-known (see [27, Lemma 10.3]), that the mapping $\varphi:({\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}})\times{\boldsymbol{\mathit{B}}}\rightarrow{\boldsymbol{\mathit{A}}}$ defined as $\langle{x,y}\rangle\mapsto p(x)+\iota(y)$ is a group isomorphism. We want to show that $\varphi$ is an $\ell$-group isomorphism between $({\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}})\overrightarrow{\times}{\boldsymbol{\mathit{B}}}$ and $\boldsymbol{\mathit{A}}$} as well. We will show that $\varphi$ is order preserving. Let us pick $\langle{a,b}\rangle,\langle{c,d}\rangle\in(A/B)\overrightarrow{\times}B$ such that $\langle{a,b}\rangle\leq\langle{c,d}\rangle$. We distinguish two cases:
  1. 1.
    If $a<c$ we get $\pi(p(a)+\iota(b))<\pi(p(c)+\iota(d))$ since
    $\pi(p(a)+\iota(b))=\pi(p(a))+\pi(\iota(b))=a<c=\pi(p(c))+\pi(\iota(d))=\pi(p(c)+\iota(d)).$
    Since $\pi$ is order preserving and $\boldsymbol{\mathit{A}}$} is totally ordered we derive from $\pi(p(a)+\iota(b))<\pi(p(c)+\iota(d))$ that $p(a)+\iota(b)<p(c)+\iota(d)$.
  2. 2.
    If $a=c$ then $b\leq d$ so clearly $p(a)+\iota(b)\leq p(c)+\iota(d)$.
This proves that $\varphi$ is an $\ell$-group homomorphism. It remains to prove that $\varphi^{-1}$ preserves ordering as well. This follows easily, from the fact that $\boldsymbol{\mathit{A}}$} is a totally ordered algebra and $\varphi$ is an order preserving isomorphism. Thus $\varphi$ is $\ell$-group isomorphism between $({\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}})\overrightarrow{\times}{\boldsymbol{\mathit{B}}}$ and $\boldsymbol{\mathit{A}}$}. ∎
One can easily prove a pointed version of this lemma, by just verifying, that the isomorphism in the proof of the Lemma 3.2 preserves the point structure.
Lemma 3.3.
Let ${\boldsymbol{\mathit{A}}}_{b}$ be a finitely generated totally ordered pointed Abelian $\ell$-group and ${\boldsymbol{\mathit{B}}}_{b}$ be a convex pointed $\ell$-subgroup of ${\boldsymbol{\mathit{A}}}_{b}$. Then ${\boldsymbol{\mathit{A}}}_{b}\cong({\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}})_{0}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}_{b}$.
Definition 3.4.
Let ${\boldsymbol{\mathit{A}}}_{a}\in p\mathbb{AL}$. We say $a$ is a strong unit of $\boldsymbol{\mathit{A}}$} if for each $b\in A$ there is $z\in\mathbb{Z}$ such that $z\cdot a\geq b$. We say that a pointed Abelian $\ell$-group ${\boldsymbol{\mathit{A}}}_{a}$ is strongly pointed, whenever $a$ is a strong unit of $\boldsymbol{\mathit{A}}$}.
Lemma 3.5.
Let ${\boldsymbol{\mathit{A}}}_{a}$ be a non-trivial pointed Abelian $\ell$-group. Moreover, assume there exists a strong unit of $\boldsymbol{\mathit{A}}$}. Then ${\boldsymbol{\mathit{Z}}}_{0}\overrightarrow{\times}{\boldsymbol{\mathit{A}}}_{a}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{A}}}_{a})$.
Proof.
Let us denote a strong unit of $\boldsymbol{\mathit{A}}$} by $b$. Consider an ultrapower of ${\boldsymbol{\mathit{A}}}_{a}$ defined as ${\boldsymbol{\mathit{C}}}=\prod_{i\in\omega}{\boldsymbol{\mathit{A}}}_{a}/\mathfrak{U}$, where $\mathfrak{U}$ is a non-principal ultrafilter on $\omega$ and let $\iota:d\mapsto\langle{d,\dots,d}\rangle$ be the canonical embedding of $\boldsymbol{\mathit{A}}$} into $\boldsymbol{\mathit{C}}$}. Consider the element $c\in C$ defined as $c=\langle{b,2b,3b,\dots}\rangle$. Clearly, $c>\iota(d)$ for any $d\in{\boldsymbol{\mathit{A}}}$. Let us denote by $\boldsymbol{\mathit{B}}$} the $\ell$-subgroup of $\boldsymbol{\mathit{C}}$} generated by $\iota[A]\cup\{c\}$. For any $d\in A$ we have $c<\iota(d)<c$ thus we can uniquely express any element of $B$ as $n\cdot c+\iota(d)$ for some $n\in\mathbb{Z}$ and $d\in A$. For any $n_{1},n_{2}\in\mathbb{Z}$ and $d_{1},d_{2}\in A$ we have $n_{1}\cdot c+\iota(d_{1})\geq n_{2}\cdot c+\iota(d_{2})$ iff $n_{1}>n_{2}$ or $n_{1}=n_{2}$ and $d_{1}\geq d_{2}$. This proves ${\boldsymbol{\mathit{B}}}_{\iota(a)}\cong{\boldsymbol{\mathit{Z}}}_{0}\overrightarrow{\times}{\boldsymbol{\mathit{A}}}_{a}$. Since ${\boldsymbol{\mathit{B}}}_{\iota(a)}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{A}}}_{a})$, we conclude ${\boldsymbol{\mathit{Z}}}_{0}\overrightarrow{\times}{\boldsymbol{\mathit{A}}}_{a}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{A}}}_{a})$. ∎
Corollary 3.6.
$p\mathbb{AL}^{0}$ is the smallest nontrivial subvariety of $p\mathbb{AL}$. Alternatively, we can say that any non-trivial proper subvariety of $p\mathbb{AL}$ contains $p\mathbb{AL}^{0}$ as a subvariety.
Proof.
Let $\mathbb{K}$}} be a nontrivial subvariety of $p\mathbb{AL}$. Take any ${\boldsymbol{\mathit{A}}}_{a}\in{{\mathbb{K}}}$. Take the subalgebra of ${\boldsymbol{\mathit{A}}}_{a}$ generated by $a$. Such an algebra has to be isomorphic to ${\boldsymbol{\mathit{Z}}}_{1}$, ${\boldsymbol{\mathit{Z}}}_{0}$ or ${\boldsymbol{\mathit{Z}}}_{-1}$. If we get ${\boldsymbol{\mathit{Z}}}_{1}\in{{\mathbb{K}}}$ or ${\boldsymbol{\mathit{Z}}}_{-1}\in{{\mathbb{K}}}$, by Lemma 3.5 we get ${\boldsymbol{\mathit{Z}}}_{0}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{1}\in{{\mathbb{K}}}$ or ${\boldsymbol{\mathit{Z}}}_{0}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{-1}\in{{\mathbb{K}}}$. Since $\mathbb{K}$}} is closed under $\mathbf{H}$, we get ${\boldsymbol{\mathit{Z}}}_{0}\in{{\mathbb{K}}}$ in all cases. Therefore, ${\boldsymbol{\mathit{Z}}}_{0}\in{{\mathbb{K}}}$ and by Lemma 2.11 we conclude that $p\mathbb{AL}^{0}\subseteq{{\mathbb{K}}}$. ∎
Using this corollary we get the stronger version of Lemma 2.11.
Lemma 3.7.
  1. 1.
    $p\mathbb{AL}=\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{-1},{\boldsymbol{\mathit{R}}}_{1})$.
  2. 2.
    $p\mathbb{AL}^{+}=\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{1})$.
  3. 3.
    $p\mathbb{AL}^{-}=\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{-1})$.
  4. 4.
    $p\mathbb{AL}^{0}=\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{0})=\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{0})$.
We will need to understand how partial embedding interacts with lexicographic products.
Lemma 3.8.
Let ${\boldsymbol{\mathit{A}}},{\boldsymbol{\mathit{B}}},{\boldsymbol{\mathit{C}}},{\boldsymbol{\mathit{D}}}$ be $\ell$-group chains and assume $\boldsymbol{\mathit{A}}$} partially embeds into $\boldsymbol{\mathit{B}}$}. Then ${\boldsymbol{\mathit{C}}}\overrightarrow{\times}{\boldsymbol{\mathit{A}}}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}$ partially embeds into ${\boldsymbol{\mathit{C}}}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}$.
Equivalently, ${\boldsymbol{\mathit{A}}}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{B}}})$ implies ${\boldsymbol{\mathit{C}}}\overrightarrow{\times}{\boldsymbol{\mathit{A}}}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{C}}}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}\overrightarrow{\times}{\boldsymbol{\mathit{D}}})$.
Proof.
Let $\{\varphi_{F}\}_{F\subseteq A,|F|<\omega}$ be a family of partial embeddings from $\boldsymbol{\mathit{A}}$} to $\boldsymbol{\mathit{B}}$}. Denote by $\pi_{A}$ the projection from ${\boldsymbol{\mathit{C}}}\overrightarrow{\times}{\boldsymbol{\mathit{A}}}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}$ to $\boldsymbol{\mathit{A}}$}. We claim that $\{{\psi_{G}\}}_{G\subseteq C\times A\times D,|G|<\omega}$, where $\psi_{G}:\langle{c,a,d}\rangle\mapsto\langle{c,\varphi_{\pi_{A}[G]}(a),d}\rangle$, is a family of partial embeddings from ${\boldsymbol{\mathit{C}}}\overrightarrow{\times}{\boldsymbol{\mathit{A}}}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}$ to ${\boldsymbol{\mathit{C}}}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}$.
Let us fix $G$ a finite subset of $C\times A\times D$. We show $\psi_{G}$ is a partial embedding. Assume $\langle{c_{1},a_{1},d_{1}}\rangle,\langle{c_{2},a_{2},d_{2}}\rangle,\langle{c_{1}+c_{2},a_{1}+a_{2},d_{1}+d_{2}}\rangle\in G$. Consequently $a_{1},a_{2},a_{1}+a_{2}\in\pi_{A}[G]$ and we have
$\psi_{G}(\langle{c_{1},a_{1},d_{1}}\rangle)+\psi_{G}(\langle{c_{2},a_{2},d_{2}}\rangle)=\langle{c_{1},\varphi_{\pi_{A}[G]}(a_{1}),d_{1}}\rangle+\langle{c_{2},\varphi_{\pi_{A}[G]}(a_{2}),d_{2}}\rangle=$
$\langle{c_{1}+c_{2},\varphi_{\pi_{A}[G]}(a_{1})+\varphi_{\pi_{A}[G]}(a_{2}),d_{1}+d_{2}}\rangle=\langle{c_{1}+c_{2},\varphi_{\pi_{A}[G]}(a_{1}+a_{2}),d_{1}+d_{2}}\rangle=$
$=\psi_{G}(\langle{c_{1}+c_{2},a_{1}+a_{2},d_{1}+d_{2}}\rangle).$
In a similar fashion we can check that if $\langle{c,a,d}\rangle,\langle{-c,-a,-d}\rangle\in G$ we have $\psi_{G}\langle{c,a,d}\rangle=\psi_{G}\langle{-c,-a,-d}\rangle$ and always we have $\psi_{G}\langle{0,0,0}\rangle=\langle{0,0,0}\rangle$.
Therefore, $\psi_{G}$ preserves addition, subtraction and zero constant. It remains to check $\psi_{G}$ preserves operations $\lor,\land$. Since ${\boldsymbol{\mathit{A}}},{\boldsymbol{\mathit{B}}},{\boldsymbol{\mathit{C}}},{\boldsymbol{\mathit{D}}}$ are chains it is enough to show that $\psi_{G}$ preserves the lattice ordering. Let us assume that $\langle{c_{1},a_{1},d_{1}}\rangle,\langle{c_{2},a_{2},d_{2}}\rangle\in G$ and $\langle{c_{1},a_{1},d_{1}}\rangle<\langle{c_{2},a_{2},d_{2}}\rangle$. Let us distinguish three cases:
  1. 1.
    If $c_{1}<c_{2}$ we get
    $\psi_{G}(\langle{c_{1},a_{1},d_{1}}\rangle)=\langle{c_{1},\varphi_{\pi_{A}[G]}(a_{1}),d_{1}}\rangle<\langle{c_{2},\varphi_{\pi_{A}[G]}(a_{2}),d_{2}}\rangle=\psi_{G}(\langle{c_{2},a_{2},d_{2}}\rangle).$
  2. 2.
    If $c_{1}=c_{2}$ and $a_{1}<a_{2}$ we get $\varphi_{\pi_{A}[G]}(a_{1})<\varphi_{\pi_{A}[G]}(a_{2})$ (by injectivity and order-preservation of $\varphi_{\pi_{A}[G]}$) and thus we get
    $\psi_{G}(\langle{c_{1},a_{1},d_{1}}\rangle)=\langle{c_{1},\varphi_{\pi_{A}[G]}(a_{1}),d_{1}}\rangle<\langle{c_{1},\varphi_{\pi_{A}[G]}(a_{2}),d_{2}}\rangle=\psi_{G}(\langle{c_{2},a_{2},d_{2}}\rangle).$
  3. 3.
    If $c_{1}=c_{2},a_{1}=a_{2}$ and $d_{1}<d_{2}$ we get
    $\psi_{G}(\langle{c_{1},a_{1},d_{1}}\rangle)=\langle{c_{1},\varphi_{\pi_{A}[G]}(a_{1}),d_{1}}\rangle<\langle{c_{1},\varphi_{\pi_{A}[G]}(a_{1}),d_{2}}\rangle=\psi_{G}(\langle{c_{2},a_{2},d_{2}}\rangle).$
This shows $\psi_{G}$ is indeed a partial embedding of $G$ into ${\boldsymbol{\mathit{C}}}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}$. Since the finite set $G$ was arbitrary, we get that ${\boldsymbol{\mathit{C}}}\overrightarrow{\times}{\boldsymbol{\mathit{A}}}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}$ partially embeds into ${\boldsymbol{\mathit{C}}}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}$. ∎
We can easily observe that the lemma also holds for pointed $\ell$-groups.
Lemma 3.9.
Let ${\boldsymbol{\mathit{A}}}_{a},{\boldsymbol{\mathit{B}}}_{b},{\boldsymbol{\mathit{C}}}_{c},{\boldsymbol{\mathit{D}}}_{d}$ be pointed $\ell$-group chains and assume ${\boldsymbol{\mathit{A}}}_{a}$ partially embeds into ${\boldsymbol{\mathit{B}}}_{b}$. Then ${\boldsymbol{\mathit{C}}}_{c}\overrightarrow{\times}{\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}_{d}$ partially embeds into ${\boldsymbol{\mathit{C}}}_{c}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}_{b}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}_{d}$.
Equivalently, ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{B}}}_{b})$ implies ${\boldsymbol{\mathit{C}}}_{c}\overrightarrow{\times}{\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}_{d}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{C}}}_{c}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}_{b}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}_{d})$.
Finally, we conclude this section with the following theorem which is a strengthening of [28, Lemma 4.6].
Theorem 3.10.
Let ${\boldsymbol{\mathit{A}}}_{b}$ be a finitely generated totally ordered pointed Abelian $\ell$-group and ${\boldsymbol{\mathit{B}}}_{b}$ be its convex strongly pointed $\ell$-subgroup, with $b\neq 0$. [^4] Let us note, that that such ${\boldsymbol{\mathit{B}}}_{b}$ is the smallest pointed convex $\ell$-subgroup of ${\boldsymbol{\mathit{A}}}_{b}$. As such, ${\boldsymbol{\mathit{B}}}_{b}$ is unique as it is uniquelly determined by ${\boldsymbol{\mathit{A}}}_{b}$. Then $\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{A}}}_{b})=\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{B}}}_{b})$.
Proof.
Clearly, ${\boldsymbol{\mathit{B}}}_{b}\in\mathbf{S}({\boldsymbol{\mathit{A}}}_{b})$ and thus ${\boldsymbol{\mathit{B}}}_{b}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{A}}}_{b}).$
We prove the other implication. By Lemma 3.5 we have ${\boldsymbol{\mathit{Z}}}_{0}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}_{b}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{B}}}_{b})$. Also, by Lemma 2.11 we have $({\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}})_{0}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{0})$ and by Lemma 3.9 we have $({\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}})_{0}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}_{b}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{0}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}_{b})$. By Lemma 3.3 we get $({\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}})_{0}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}_{b}\cong{\boldsymbol{\mathit{A}}}_{b}$. Therefore we showed ${\boldsymbol{\mathit{A}}}_{b}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{B}}}_{b})$, which completes the proof. ∎
Theorem 3.10 tells us that the universal theory of any totally ordered pointed $\ell$-group is equal to the universal theory generated by its convex $\ell$-subgroup, which is strongly pointed or the universal theory generated by ${\boldsymbol{\mathit{Z}}}_{0}$. In other words, when classifying universal classes of totally ordered pointed $\ell$-groups, we can restrict our focus only on those which are strongly pointed or 0-pointed.
4 Characterization of varieties generated by a single finitely generated totally ordered $\ell$-group
In this section we describe all join-irreducible subvarieties of pointed Abelian $\ell$-groups. Although, one could obtain the result of this chapter using Theorem 3.10 from Section 3, Mundici functor and Komori classification of MV-algebras, we have chosen a different, possibly harder approach, by providing a more self-contained theory and not using the theory of MV-algebras. However, a reader familiar with Komori classification of MV-algebras will find some of the proofs here possibly familiar, since they often use similar techniques (for comparison see [7]).
Lemma 4.1.
Let ${\boldsymbol{\mathit{A}}}_{a},{\boldsymbol{\mathit{B}}}_{b},{\boldsymbol{\mathit{C}}}_{c}$ be totally ordered pointed Abelian $\ell$-groups and ${\boldsymbol{\mathit{D}}}_{d}$ be a pointed Abelian $\ell$-group. Let $\psi:{\boldsymbol{\mathit{A}}}_{a}\rightarrow{\boldsymbol{\mathit{B}}}_{b}$ be an injective homomorphism. Then $\varphi:{\boldsymbol{\mathit{C}}}_{c}\overrightarrow{\times}{\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}_{d}\rightarrow{\boldsymbol{\mathit{C}}}_{c}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}_{b}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}_{d}$ defined as $\langle{x,y,z}\rangle\mapsto\langle{x,\psi(y),z}\rangle$ is an injective homomorphism as well.
Moreover, if $\psi$ is an isomorphism then also $\varphi$ is an isomorphism.
Proof.
First, $\langle{x,y,z}\rangle\mapsto\langle{x,\psi(y),z}\rangle$ is clearly an injective homomorphism of groups ${\boldsymbol{\mathit{C}}}_{c}\times{\boldsymbol{\mathit{A}}}_{a}\times{\boldsymbol{\mathit{D}}}_{d}$ and ${\boldsymbol{\mathit{C}}}_{c}\times{\boldsymbol{\mathit{B}}}_{b}\times{\boldsymbol{\mathit{D}}}_{d}$.
We check that $\varphi$ preserves the lattice operations of lexicographic product.
For each $x_{1},x_{2}\in C,y_{1},y_{2}\in A$ and $z_{1},z_{2}\in D$ we have
$\varphi(\langle{x_{1},y_{1},z_{1}}\rangle)\lor\varphi(\langle{x_{2},y_{2},z_{2}}\rangle)=\langle{x_{1},\psi(y_{1}),z_{1}}\rangle\lor\langle{x_{2},\psi(y_{2}),z_{2}}\rangle=$
[…]
$\varphi(\langle{x_{1},y_{1},z_{1}}\rangle\lor\langle{x_{2},y_{2},z_{2}}\rangle).$
This shows $\varphi$ preserves $\lor$. Similarly we can show that $\varphi$ preserves $\land$. Therefore, $\varphi$ is an injective homomorphism.
In the case, when $\psi$ is an isomorphism, there exists the inverse isomorphism $\psi^{-1}:{\boldsymbol{\mathit{B}}}_{b}\rightarrow{\boldsymbol{\mathit{A}}}_{a}$. By the previous part of the proof $\varphi^{-1}:{\boldsymbol{\mathit{C}}}_{c}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}_{b}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}_{d}\rightarrow{\boldsymbol{\mathit{C}}}_{c}\overrightarrow{\times}{\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}_{d}$ defined as $\langle{x,y,z}\rangle\mapsto\langle{x,\psi^{-1}(y),z}\rangle$ is an injective homomorphism. Clearly, $\varphi^{-1}$ is also inverse to $\varphi$, which shows $\varphi$ has to be an isomorphism. ∎
Lemma 4.2.
Let $a_{1},\dots,a_{m}\in\mathbb{R}$ and let $f:({\boldsymbol{\mathit{R}}}\overrightarrow{\times}{\boldsymbol{\mathit{R}}})^{m}\rightarrow{\boldsymbol{\mathit{R}}}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}$ be a pointed $\ell$-group term function. Let $\pi_{2}:{\boldsymbol{\mathit{R}}}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}\rightarrow{\boldsymbol{\mathit{R}}}$ be a projection to the second coordinate. Then the $m$-ary function
$g(x_{1},\dots,x_{m}):=\pi_{2}(f(\langle{a_{1},x_{1}}\rangle,\dots,\langle{a_{m},x_{m}}\rangle))$
is a continuous function from ${\boldsymbol{\mathit{R}}}^{m}$ to $\boldsymbol{\mathit{R}}$}. In particular, every pointed $\ell$-group term function on $\boldsymbol{\mathit{R}}$} is continuous.
Proof.
We show that $g$ is a composition of continuous functions. Clearly, constants are continuous functions. Moreover, addition and subtraction are defined component-wise thus they are continuous functions.
It remains to check maximum and minimum. For each constants $b_{1},b_{2}\in\mathbb{R}$ we define
[…]
and
$x_{1}\land_{b_{1},b_{2}}x_{2}=\begin{cases}x_{2}&\text{if }b_{1}>b_{2}\\ x_{1}&\text{if }b_{1}<b_{2}\\ x_{1}\land x_{2}&\text{if }b_{1}=b_{2}\end{cases}.$
Thus for all $b_{1},b_{2}\in\mathbb{R}$ we obtain that $\lor_{b_{1},b_{2}}$ and $\land_{b_{1},b_{2}}$ are continuous functions. Since $g$ is a composition of addition, subtraction, constants and functions $\lor_{b_{i},b_{j}}$ and $\land_{b_{i},b_{j}}$ for some $b_{i},b_{i}\in\mathbb{R}$, we obtain that $g$ is continuous as well. ∎
The following lemma is the generalization of [7, Propositions 8.1.1 and 8.1.2]
Lemma 4.3.
  1. 1.
    Let ${\boldsymbol{\mathit{A}}}_{a}$ be a pointed $\ell$-subgroup of ${\boldsymbol{\mathit{R}}}_{a}$ and $A$ be a dense subset of $\mathbb{R}$. Then $\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{a})=\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a})$.
  2. 2.
    Let $I$ be an infinite set of positive integers, $b>0$ and ${\boldsymbol{\mathit{A}}}_{a}$ be a pointed $\ell$-subgroup of ${\boldsymbol{\mathit{R}}}_{a}$. Then $\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{b})=\mathbf{HSP}(\{{\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{n}\mid n\in I\})$.
  3. 3.
    Let $I$ be an infinite set of negative integers, $b<0$ and ${\boldsymbol{\mathit{A}}}_{a}$ be a pointed $\ell$-subgroup of ${\boldsymbol{\mathit{R}}}_{a}$. Then $\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{b})=\mathbf{HSP}(\{{\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{n}\mid n\in I\})$.
Proof.
  1. 1.
    Clearly, ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{a})$.
    For the other inclusion assume that some equation $\alpha(\vec{x})=0$ does not hold in ${\boldsymbol{\mathit{R}}}_{a}$. Thus there is $\vec{r}\in\mathbb{R}^{m}$, where $m$ is the number of variables in $\alpha$, such that $\alpha(\vec{r})\neq 0$. By Lemma 4.2 the function $\vec{x}\mapsto\alpha(\vec{x})$ is continuous. Since the set $\{x\in\mathbb{R}\ \mid x\neq 0\}$ is open in $\mathbb{R}$, the set $O:=\{\vec{x}\in\mathbb{R}^{m}\mid\alpha(\vec{x})\neq 0\}$ is open in $\mathbb{R}^{m}$. We know that $\vec{r}\in O$ thus the set $O$ is nonempty. Since $A$ is a dense subset of $\mathbb{R}$, by [26, Theorem 19.5] $A^{m}$ is a dense subset of $\mathbb{R}^{m}$ and therefore $A^{m}\cap O\neq\emptyset$. Thus there exist $\vec{a}\in A^{m}\cap O$ such that $\alpha(\vec{a})\neq 0$. This shows that the equation $\alpha(\vec{x})=0$ is not valid in ${\boldsymbol{\mathit{A}}}_{a}$. Since $\alpha(\vec{x})=0$ was an arbitrary equation we derive that ${\boldsymbol{\mathit{R}}}_{a}\in\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a})$. Thus $\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a})=\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{a})$.
  2. 2.
    Without loss of generality we can assume $b=1$, since ${\boldsymbol{\mathit{R}}}_{1}\cong{\boldsymbol{\mathit{R}}}_{b}$ for any $b>0$ and by Lemma 4.1 ${\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{1}\cong{\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{b}$ for any such $b$. Since ${\boldsymbol{\mathit{Z}}}_{n}\in\mathbf{IS}({\boldsymbol{\mathit{R}}}_{1})$ for each $n\in I$, by Lemma 4.1 we get ${\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{n}\in\mathbf{IS}({\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{1})$ for each $n\in I$.
    For the other inclusion assume that some equation $\alpha(\langle{\vec{y},\vec{x}}\rangle)=0$ does not hold in ${\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{1}$. Let $m$ be the number of variables in $\alpha$. There exist $\langle{\vec{s},\vec{r}}\rangle\in A^{m}\times\mathbb{R}^{m}$ such that $\alpha(\langle{\vec{s},\vec{r}}\rangle)\neq 0$ in ${\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{1}$. By Lemma 4.2 the function $\vec{x}\rightarrow\alpha(\vec{s},\vec{x})$ is continuous and thus the set $O:=\{\vec{x}\in{\boldsymbol{\mathit{R}}}^{m}\mid\alpha(\langle{\vec{s},\vec{x}}\rangle)\neq 0\}$ is open in ${\boldsymbol{\mathit{R}}}^{m}$ and contains $\vec{r}$. Therefore, there is $\epsilon\in\mathbb{R}$ such that $V:=\{\vec{x}\mid|\vec{x}-\vec{r}|<m\cdot\epsilon\}\subseteq O$. We fix $n\in I$ such that $n>\frac{\sqrt{m}}{2\epsilon}$. Let us consider $\ell$-group ${{\boldsymbol{\mathit{Z}}}[\frac{1}{n}]}_{1}$. We have $\epsilon>\frac{\sqrt{m}}{2n}$ and by [10, Section 5, Chapter 4] it follows that ${{\boldsymbol{\mathit{Z}}}[\frac{1}{n}]}_{1}\cap V\neq\emptyset$. Since $V\subseteq O$ the equation $\alpha(\vec{s},\vec{x})=0$ is not valid in ${\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{{\boldsymbol{\mathit{Z}}}[\frac{1}{n}]}_{1}$. By Lemma 4.1 we get ${\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{{\boldsymbol{\mathit{Z}}}[\frac{1}{n}]}_{1}\cong{\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{n}$, which proves ${\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{1}\in\mathbf{HSP}(\{{\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{n}\mid n\in I\})$.
  3. 3.
    This can be proved similarly as the previous point.
Lemma 4.4.
Let $a,k\in\mathbb{Z}$. Then ${\boldsymbol{\mathit{Z}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{k}\cong{\boldsymbol{\mathit{Z}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{k+a}$.
Proof.
Consider the mapping $\varphi:{\boldsymbol{\mathit{Z}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{k}\rightarrow{\boldsymbol{\mathit{Z}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{k+a}$ defined as $\langle{x,y}\rangle\mapsto\langle{x,y+x}\rangle$. We show $\varphi$ is an isomorphism of $\ell$-groups. Clearly, this is a group isomorphism with inverse mapping $\varphi^{-1}$ defined as $\langle{x,y}\rangle\mapsto\langle{x,y-x}\rangle$. It remains to check $\varphi$ and $\varphi^{-1}$ are order preserving. Assume $\langle{a_{1},b_{1}}\rangle\leq\langle{a_{2},b_{2}}\rangle$ for some $a_{1},a_{2},b_{1},b_{2}\in\mathbb{Z}$.
  1. 1.
    If $a_{1}<a_{2}$ we have
    $\varphi(\langle{a_{1},b_{1}}\rangle)=\langle{a_{1},a_{1}+b_{1}}\rangle<\langle{a_{2},a_{2}+b_{2}}\rangle=\varphi(\langle{a_{2},b_{2}}\rangle).$
  2. 2.
    If $a_{1}=a_{2}$ we obtain $b_{1}\leq b_{2}$ and $a_{1}+b_{1}\leq a_{2}+b_{2}$. Therefore,
    $\varphi(\langle{a_{1},b_{1}}\rangle)=\langle{a_{1},a_{1}+b_{1}}\rangle\leq\langle{a_{1},a_{2}+b_{2}}\rangle=\langle{a_{2},a_{2}+b_{2}}\rangle=\varphi(\langle{a_{2},b_{2}}\rangle).$
This shows that $\varphi$ preserves the ordering. In similar way one can show that $\varphi^{-1}$ preserves the ordering as well. Thus $\varphi$ is indeed an isomorphism. ∎
Definition 4.5.
Let ${\boldsymbol{\mathit{A}}}_{a}$ be a totally ordered pointed Abelian $\ell$-group and $a\neq 0$. We call the unique maximal convex ideal $I$ of ${\boldsymbol{\mathit{A}}}_{a}$ such that $a\notin I$ a p-radical, and we denote it by $\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$. Furthermore, we say that ${\boldsymbol{\mathit{A}}}_{a}$ is p-simple if it has a trivial p-radical, $\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})=\{0\}$.
For any Abelian pointed $\ell$-group ${\boldsymbol{\mathit{A}}}_{a}$, its radical, $\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$, is an Abelian (non-pointed) $\ell$-group. Determining a canonical point for $\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$ isn’t immediately obvious. However, when $\boldsymbol{\mathit{A}}$} is finitely generated, Lemma 3.3 provides the decomposition ${\boldsymbol{\mathit{A}}}_{a}\cong{\boldsymbol{\mathit{A}}}_{a}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\overrightarrow{\times}(\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a}))_{b}$ for a unique $b\in\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$. This means that for any finitely generated $\ell$-group, we can meaningfully consider its radical as a pointed Abelian $\ell$-group. We will define its point, $\mathtt{f}^{\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})}$, such that the decomposition ${\boldsymbol{\mathit{A}}}_{a}\cong{\boldsymbol{\mathit{A}}}_{a}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\overrightarrow{\times}\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$ holds. The preceding decomposition makes it possible to treat $\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$ as a pointed Abelian $\ell$-group. Henceforth, we adopt the convention that when we do so, its point is the one uniquely determined by this decomposition.
Let us note that here the definition of p-simple and p-radical does not coincide with the classical universal algebraic meaning of an $\ell$-group being simple and the notion of the radical of $\ell$-group. When we say that a pointed Abelian $\ell$-group ${\boldsymbol{\mathit{A}}}_{a}$ is p-simple or $\boldsymbol{\mathit{C}}$} is a p-radical of ${\boldsymbol{\mathit{A}}}_{a}$, we are saying something much weaker than an Abelian $\ell$-group $\boldsymbol{\mathit{A}}$} being simple or that $\boldsymbol{\mathit{C}}$} is a radical of $\boldsymbol{\mathit{A}}$}. One can show that the notion of p-radical and radical coincide (as well the notions of simple and p-simple) in the case of $\ell$-groups with a strong unit.
This definition gives us the following useful property: Let ${\boldsymbol{\mathit{A}}}_{a}$ be a totally ordered strongly pointed Abelian $\ell$-group and let ${\boldsymbol{\mathit{B}}}_{a}$ be its totally ordered pointed Abelian $\ell$-subgroup and $a\neq 0$. It can be shown that $\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})=\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{B}}}_{a})$. Since ${\boldsymbol{\mathit{B}}}_{a}$ is strongly pointed, the quotient ${\boldsymbol{\mathit{B}}}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{B}}}_{a})$ has to be a simple $\ell$-group and hence, according to Hölder’s Theorem (see Theorem 2.8), ${\boldsymbol{\mathit{B}}}_{a}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{B}}}_{a})$ is isomorphic to a pointed $\ell$-subgroup of the real numbers. Since any pointed Abelian $\ell$-group ${\boldsymbol{\mathit{A}}}_{a}$ with $a\neq 0$ has a unique convex strongly pointed $\ell$-subgroup, we can state the following definition.
Definition 4.6.
Let ${\boldsymbol{\mathit{A}}}_{a}$ be a totally ordered pointed Abelian $\ell$-group and ${\boldsymbol{\mathit{B}}}_{a}$ be its convex strongly pointed $\ell$-subgroup. We define rank of ${\boldsymbol{\mathit{A}}}_{a}$ as
[…]
Lemma 4.7.
Let ${\boldsymbol{\mathit{A}}}_{a}$ be a totally ordered p-simple pointed Abelian $\ell$-group and $\mathbf{rank}({\boldsymbol{\mathit{A}}}_{a})=n\in\mathbb{Z}\setminus\{0\}$. Then ${\boldsymbol{\mathit{A}}}_{a}\cong{\boldsymbol{\mathit{Z}}}_{n}$.
Proof.
Follows from the definitions of rank and p-simple $\ell$-group. ∎
Lemma 4.8.
Let ${\boldsymbol{\mathit{A}}}_{a}$ be a finitely generated totally ordered non-p-simple Abelian $\ell$-group with rank $n$. Then $\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a})=\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$.
Proof.
By Theorem 3.10 we can assume without loss of generality that ${\boldsymbol{\mathit{A}}}_{a}$ is strongly pointed. Since $\mathbf{rank}({\boldsymbol{\mathit{A}}}_{a})=n$, by Lemma 3.2 and Lemma 4.1 we obtain ${\boldsymbol{\mathit{A}}}_{a}\cong{\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$. We first show ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\in\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a}))$.
Since ${\boldsymbol{\mathit{A}}}_{a}$ is finitely generated, $\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$ has a strong unit and thus Lemma 3.5 gives us ${\boldsymbol{\mathit{Z}}}_{0}\overrightarrow{\times}\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{ISP}_{\mathrm{U}}(\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a}))$ and thus by Lemma 3.9 we obtain ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\overrightarrow{\times}\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a}))$. Finally, ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\in\mathbf{H}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\overrightarrow{\times}\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a}))$ and thus ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\in\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a}))$.
We have to show the other inclusion, i.e. that ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$. By Lemma 4.4 we have ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{kn}\in\mathbf{I}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$ for each $k\in\mathbb{Z}$. Thus by Lemma 4.3 we obtain ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{1},{\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{-1}\in\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$. Also, by Lemma 2.11 we have ${\boldsymbol{\mathit{R}}}_{0}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{0})$ and thus by Lemma 3.9 and Lemma 4.1 we have ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{0}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$.
Since $\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$ is totally ordered, by Lemma 2.11 we get $\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{-1},{\boldsymbol{\mathit{R}}}_{0},{\boldsymbol{\mathit{R}}}_{1})$, thus by [2, Theorem 5.6] we have
$\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{-1})\cup\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{0})\cup\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{1}).$
By Lemma 3.9 we obtain
${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{-1})\cup\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{0})\cup\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{1}).$
Since
${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{-1},{\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{0},{\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}_{1}\in\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}),$
we get ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}).$ This proves the claim. ∎
Lemma 4.9.
Let ${\boldsymbol{\mathit{A}}}_{a}$ be a finitely generated totally ordered Abelian $\ell$-group.
  1. 1.
    If $\mathbf{rank}({\boldsymbol{\mathit{A}}}_{a})=\infty$ then $\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a})=\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{1})$.
  2. 2.
    If $\mathbf{rank}({\boldsymbol{\mathit{A}}}_{a})=$-$\infty$ then $\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a})=\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{-1})$.
Proof.
By Theorem 3.10 we can assume without loss of generality that ${\boldsymbol{\mathit{A}}}_{a}$ is strongly pointed. Let us assume $\mathbf{rank}({\boldsymbol{\mathit{A}}}_{a})=\infty$. Since $\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$ is a maximal convex $\ell$-subgroup of ${\boldsymbol{\mathit{A}}}_{a}$, the Abelian $\ell$-group ${\boldsymbol{\mathit{A}}}_{a}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$ has to be p-simple and thus by Theorem 2.8 ${\boldsymbol{\mathit{A}}}_{a}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{IS}({\boldsymbol{\mathit{R}}}_{b})$ for some $0<b\in\mathbb{R}$. Let us without loss of generality assume ${\boldsymbol{\mathit{A}}}_{a}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{S}({\boldsymbol{\mathit{R}}}_{1})$. By Lemma 2.9 we obtain ${\boldsymbol{\mathit{A}}}_{a}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\cong{\boldsymbol{\mathit{Z}}}_{n}$ for some $n\in\mathbb{Z}$ or the universe of ${\boldsymbol{\mathit{A}}}_{a}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$ is a dense subset of $\mathbb{R}$. Since $\mathbf{rank}({\boldsymbol{\mathit{A}}}_{a})=\infty$, the universe of ${\boldsymbol{\mathit{A}}}_{a}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$ is dense in $\mathbb{R}$. Thus by Lemma 4.3 we get $\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a}))=\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{1})$. Since ${\boldsymbol{\mathit{A}}}_{a}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{H}({\boldsymbol{\mathit{A}}}_{a})$ and by Lemma 3.7 ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{1})$ we obtain $\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{1})=\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a})$.
The case $\mathbf{rank}({\boldsymbol{\mathit{A}}}_{a})=$-$\infty$ is analogous. ∎
All together we obtained the following characterization.
Theorem 4.10.
Let ${\boldsymbol{\mathit{A}}}_{a}$ be a finitely generated totally ordered pointed Abelian $\ell$-group. Then the following holds.
  1. 1.
    $\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a})=\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{0})=\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{0})$ iff $a=0$.
  2. 2.
    $\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a})=\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n})$ iff ${\boldsymbol{\mathit{A}}}_{a}$ is p-simple and $\mathbf{rank}({\boldsymbol{\mathit{A}}}_{a})=n$.
  3. 3.
    $\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a})=\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$ iff ${\boldsymbol{\mathit{A}}}_{a}$ is non-p-simple and $\mathbf{rank}({\boldsymbol{\mathit{A}}}_{a})=n$.
  4. 4.
    $\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a})=\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{1})$ iff $\mathbf{rank}({\boldsymbol{\mathit{A}}}_{a})=\infty$.
  5. 5.
    $\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a})=\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{-1})$ iff $\mathbf{rank}({\boldsymbol{\mathit{A}}}_{a})=$-$\infty$.
5 Axiomatization of subvarieties of $p\mathbb{AL}$
In this section we will introduce several equations which we later use for axiomatizations of subvarieties of $p\mathbb{AL}$. We will introduce three important equations. It should be noted that all our equations are valid in all negatively pointed $\ell$-groups.
Lemma 5.1.
Let ${\boldsymbol{\mathit{A}}}_{a}$ be a totally ordered pointed Abelian $\ell$-group and $n\geq 0$. Let us consider the following equation.
$(n\cdot x-\mathtt{f})\lor(-x)\geq 0.$ (s-rankn)
This equation is satisfied in ${\boldsymbol{\mathit{A}}}_{a}$ iff $a\leq 0$ or ${\boldsymbol{\mathit{A}}}_{a}$ is p-simple with $\mathbf{rank}({\boldsymbol{\mathit{A}}}_{a})\leq n$.
Proof.
By Theorem 3.10 we can assume ${\boldsymbol{\mathit{A}}}_{a}$ is strongly pointed. Thus by Lemma 4.7 we need to check that
${\boldsymbol{\mathit{A}}}_{a}\vDash\mathrm{(\text{s-rank}_{n})}\text{ iff }a\leq 0\text{ or }{\boldsymbol{\mathit{A}}}\cong{\boldsymbol{\mathit{Z}}}_{m}\text{ for some }m\leq n.$
Let $e$ be a fixed evaluation on ${\boldsymbol{\mathit{A}}}_{a}$. We have ${\boldsymbol{\mathit{A}}}_{a}\vDash_{e}\mathrm{(\text{s-rank}_{n})}$ iff ${\boldsymbol{\mathit{A}}}_{a}\vDash_{e}e(x)\leq 0$ or ${\boldsymbol{\mathit{A}}}_{a}\vDash_{e}n\cdot e(x)\geq a$. Therefore ${\boldsymbol{\mathit{A}}}_{a}\vDash\mathrm{(\text{s-rank}_{n})}$ iff ${\boldsymbol{\mathit{A}}}_{a}\vDash n\cdot e(x)\geq a$ for all evaluations $e$, such that $e(x)>0$.
We need to distinguish several cases.
  1. 1.
    For $a\leq 0$ and $e(x)>0$ we have $n\cdot e(x)>0\geq a$. Thus ${\boldsymbol{\mathit{A}}}_{a}\vDash\mathrm{(\text{s-rank}_{n})}$ for $a\leq 0$.
  2. 2.
    For ${\boldsymbol{\mathit{A}}}_{a}={\boldsymbol{\mathit{Z}}}_{m}$, $0<m\leq n$ and $0<e(x)$ we have $n\cdot e(x)\geq n\cdot 1\geq m$. Therefore ${\boldsymbol{\mathit{Z}}}_{m}\vDash\mathrm{(\text{s-rank}_{n})}$ for $m\leq n$.
  3. 3.
    For ${\boldsymbol{\mathit{A}}}_{a}={\boldsymbol{\mathit{Z}}}_{m}$, $0\leq n<m$ set $e(x)=1$. We have $n\cdot e(x)=n<m$, thus ${\boldsymbol{\mathit{Z}}}_{m}\nvDash\mathrm{(\text{s-rank}_{n})}$ for $m>n$.
  4. 4.
    For $a>0$ and ${\boldsymbol{\mathit{A}}}_{a}$ non-p-simple we have $\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\neq 0$ and by Lemma 3.2 there is an isomorphism $\iota:{\boldsymbol{\mathit{A}}}_{a}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\overrightarrow{\times}\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\rightarrow{\boldsymbol{\mathit{A}}}_{a}$. Pick $0<v\in\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$ and we consider evaluation $e$ on $\boldsymbol{\mathit{A}}$}, where $e(x)=\iota(\langle{0,v}\rangle)$. Since $n\cdot v\in\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$, we have $n\cdot e(x)=\iota(\langle{0,n\cdot v}\rangle)<a$, thus showing ${\boldsymbol{\mathit{A}}}_{a}\nvDash\mathrm{(\text{s-rank}_{n})}$ for non-p-simple ${\boldsymbol{\mathit{A}}}_{a}$ with $a>0$.
Let us note that $\mathrm{(\text{s-rank}_{0})}$ is equivalent to the equation $f\leq 0$. For $n\in\mathbb{Z}$ we denote $\operatorname{div}(z)$ to be a set of all divisors of $z$.
Lemma 5.2.
Let $m\in\mathbb{Z}$, $n,p\in\mathbb{N}$ and $m\leq n$.
Let us consider the following equation.
$((n+1)\cdot((p\cdot x-\mathtt{f})\lor(\mathtt{f}-p\cdot x))-\mathtt{f})\lor-x\geq 0$ (divp,n)
The following conditions are equivalent.
  1. 1.
    ${\boldsymbol{\mathit{Z}}}_{m}\vDash\mathrm{(div_{p,n})}$,
  2. 2.
    ${\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(div_{p,n})}$,
  3. 3.
    $m\leq 0$ or $p\notin\operatorname{div}(m)$.
Proof.
Since ${\boldsymbol{\mathit{Z}}}_{m}\in\mathbf{H}({\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$ it is enough to check that ${\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(div_{p,n})}$ if $m\leq 0$ or $p\notin\operatorname{div}(m)$ and that ${\boldsymbol{\mathit{Z}}}_{m}\nvDash\mathrm{(div_{p,n})}$ if $0<m$ and $p\in\operatorname{div}(m)$.
  1. 1.
    Let $e$ be a fixed evaluation on ${\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}$. We have
    ${\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash_{e}\mathrm{(div_{p,n})}\text{ iff }e(x)\leq 0\text{ or }(n+1)\cdot|p\cdot e(x)-\langle{m,0}\rangle|\geq\langle{m,0}\rangle.$
    Therefore ${\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(\text{div}_{p,n})}$ iff $(n+1)\cdot|p\cdot e(x)-\langle{m,0}\rangle|\geq\langle{m,0}\rangle$ for all evaluations $e$, such that $e(x)>0$. For an evaluation $e(x)>0$ and $m\leq 0$, we get
    $(n+1)\cdot|p\cdot e(x)-\langle{m,0}\rangle|\geq\langle{0,0}\rangle\geq\langle{m,0}\rangle.$
    This tells us that ${\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(\text{div}_{p,n})}$ for $m\leq 0$.
    Now assume $0<m\leq n,p\notin\operatorname{div}(m)$ and $e(x)=\langle{a_{1},a_{2}}\rangle$ and $a_{1}\geq 0$. Since $p\notin\operatorname{div}(m)$ we get $p\cdot a_{1}\neq m$ and thus $|p\cdot a_{1}-m|\geq 1$. Consequently,
    $(n+1)\cdot|p\cdot a_{1}-m|>m\cdot|p\cdot a_{1}-m|\geq m.$
    Thus
    $(n+1)\cdot|p\cdot\langle{a_{1},a_{2}}\rangle-\langle{m,0}\rangle|=|\langle{(n+1)\cdot(p\cdot a_{1}-m),((n+1)\cdot p\cdot a_{2})}\rangle|\geq\langle{m,0}\rangle.$
    Therefore, ${\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(div_{p,n})}$ for $m>0$ such that $p\notin\operatorname{div}(m)$.
  2. 2.
    Now assume $0<m$ and $p\mid m$. Since $p\mid m$, there is $r\in\mathbb{N}$ such that $r\cdot p=m$. Let us consider an evaluation $e$ on ${\boldsymbol{\mathit{Z}}}_{m}$, where $e(x)=r$. We have
    $(n+1)\cdot|p\cdot r-m|=(n+1)\cdot 0=0<m.$
    This shows ${\boldsymbol{\mathit{Z}}}_{m}\nvDash\mathrm{(div_{p,n})}$ if $0<m$ and $p\in\operatorname{div}(m)$.
This completes the proof. ∎
Now we have enough tools to axiomatize the variety $\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n})$.
Theorem 5.3.
Let $n\in\mathbb{N}$. The variety $\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n})$ can be axiomatized as a subvariety of $p\mathbb{AL}^{+}$ by the following set of formulas: $\{\mathrm{(\text{s-rank}_{n})}\}\cup\{\mathrm{(\text{div}_{p,{n}})}\mid p\notin\operatorname{div}(n),p<n\}$.
Proof.
By Lemma 5.1 we have ${\boldsymbol{\mathit{Z}}}_{n}\vDash\mathrm{(\text{s-rank}_{n})}$ and by Lemma 5.2 we have ${\boldsymbol{\mathit{Z}}}_{n}\vDash\mathrm{(\text{div}_{p,{n}})}$ for all $p\notin\operatorname{div}(n)$.
Now assume ${\boldsymbol{\mathit{A}}}_{a}$ is an arbitrary finitely generated subdirectly irreducible positively pointed Abelian $\ell$-group, such that ${\boldsymbol{\mathit{A}}}_{a}\vDash\mathrm{(\text{s-rank}_{n})}$ and ${\boldsymbol{\mathit{A}}}_{a}\vDash\mathrm{(\text{div}_{p,{n}})}$ for all $p\notin\operatorname{div}(n)$. We will show ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n})$.
By Theorem 2.5 ${\boldsymbol{\mathit{A}}}_{a}$ is totally ordered. By Theorem 3.10 we can assume without loss of generality that ${\boldsymbol{\mathit{A}}}_{a}$ is strongly pointed. By Lemma 5.1 we obtain ${\boldsymbol{\mathit{A}}}_{a}$ has to be p-simple with rank $n$ and thus by Lemma 4.7 ${\boldsymbol{\mathit{A}}}_{a}\cong{\boldsymbol{\mathit{Z}}}_{k}$ for some $k\leq n$. By Lemma 5.2 we obtain $k$ divides $n$. That means ${\boldsymbol{\mathit{A}}}_{a}\cong{\boldsymbol{\mathit{Z}}}_{k}\in\mathbf{IS}(Z_{n})$ and thus ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n})$. ∎
Since Theorem 5.7 also provides an axiomatization for the variety generated by ${\boldsymbol{\mathit{Z}}}_{n}$, it might seem that this would render Theorem 5.3 redundant. However, this is not the case. The varieties generated by ${\boldsymbol{\mathit{Z}}}_{n}$ are structurally simpler than the general case, as they are not generated by any $\ell$-groups of the form ${\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}$. Consequently, Theorem 5.3 provides a much simpler and more direct axiomatization than the one required by the general framework of Theorem 5.7. This is particularly appealing because the varieties generated by ${\boldsymbol{\mathit{Z}}}_{n}$ correspond via the Mundici functor to the varieties generated by single finite Łukasiewicz chains, which are among the most fundamental structures in the theory of MV-algebras.
Lemma 5.4.
Let $n\geq 0$ and $m\in\mathbb{Z}$. Let us consider the following equation.
$((2n+1)\cdot x-2\cdot\mathtt{f})\lor(\mathtt{f}-(2n+2)\cdot x)\lor-x\geq 0.$ ($rank_{n}$)
The following conditions are equivalent.
  1. 1.
    ${\boldsymbol{\mathit{Z}}}_{m}\vDash\mathrm{(\text{rank}_{n})}$,
  2. 2.
    ${\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(\text{rank}_{n})}$,
  3. 3.
    $m\leq n$.
Proof.
Since ${\boldsymbol{\mathit{Z}}}_{n}\in\mathbf{H}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$, we only have to show ${\boldsymbol{\mathit{Z}}}_{m}\nvDash\mathrm{(\text{rank}_{n})}$ for $m>n$ and ${\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(\text{rank}_{n})}$ for $m\leq n$.
  1. 1.
    First we show ${\boldsymbol{\mathit{Z}}}_{m}\nvDash\mathrm{(\text{rank}_{n})}$ for $n<m$. Since ${\boldsymbol{\mathit{Z}}}_{m}$ is Archimedean, there is a maximal $k\geq 1$ such that $k\cdot(2n+1)<2m$. Let us note, that from the maximality of $k$ it follows $2k\cdot(2n+1)\geq 2m$ and thus $k\cdot(2n+1)\geq m$. Let $e$ be an evaluation on ${\boldsymbol{\mathit{Z}}}_{m}$ such that $e(x)=k$. Now we have
    $(2n+1)\cdot e(x)-2\cdot e(\mathtt{f})=(2n+1)\cdot k-2\cdot m<0$
    by definition of $k$. Moreover, we have
    $e(\mathtt{f})-(2n+2)\cdot e(x)=m-k(2n+2)\leq k(2n+1)-k(2n+2)=-k<0.$
    Clearly, also e(x)=k<0-e(x)=-k<0. This shows ${\boldsymbol{\mathit{Z}}}_{m}\nvDash_{e}\mathrm{(\text{rank}_{n})}$ for $n<m$.
  2. 2.
    It remains to show that ${\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(\text{rank}_{n})}$ for $m\leq n$. Trivially, we have ${\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash_{e}\mathrm{(\text{rank}_{n})}$ for any evaluation $e(x)\leq 0$. Therefore, from now on we can focus only on evaluations such that $e(x)>0$. For $m\leq 0$ and evaluation $e$ such that $e(x)>0$ we have $(2n+1)\cdot e(x)-2e(\mathtt{f})\geq 0$. Thus ${\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash_{e}\mathrm{(\text{rank}_{n})}$ for $m\leq 0$.
    Now assume $m>0$. First, let us consider the case $e(x)=\langle{0,b}\rangle$ for some $b\in\mathbb{Z}$. We have
    $e(\mathtt{f})-(2n+2)\cdot e(x)=\langle{m,0}\rangle-(2n+2)\cdot\langle{0,b}\rangle=\langle{m,-(2n+2)\cdot b}\rangle>0.$
    Finally we have to check the case when $e(x)=\langle{a,b}\rangle$ for some $a,b\in\mathbb{Z}$ and $a>0$. We have
    $(2n+1)\cdot e(x)-2\cdot e(\mathtt{f})=(2n+1)\cdot\langle{a,b}\rangle-2\cdot\langle{m,0}\rangle\geq(2n+1)\cdot\langle{1,b}\rangle-2\cdot\langle{n,0}\rangle=\\ =\langle{(2n+1)-2n,(2n+1)\cdot b}\rangle=\langle{1,(2n+1)\cdot b}\rangle\geq 0.$
    This shows that indeed ${\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(\text{rank}_{n})}$ for $m\leq n$.
Now we can introduce the final equation, which we need to axiomatize the subvarieties of $\mathrm{pAb}$}. This equation is equivalent to the disjunction of the equations $\mathrm{(\text{s-rank}_{n})}$ and $\mathrm{(div_{p,n})}$.
Lemma 5.5.
Let $\boldsymbol{\mathit{A}}$} be totally ordered Abelian $\ell$-group and $0\leq p\leq n$. Let us consider the following equation.
$((n+1)\cdot((p\cdot x-\mathtt{f})\lor(\mathtt{f}-p\cdot x))-\mathtt{f})\lor(n\cdot y-\mathtt{f})\lor(-y)\geq 0$ (mixp,n)
We have ${\boldsymbol{\mathit{A}}}\vDash\mathrm{(mix_{p,n})}$ if and only if ${\boldsymbol{\mathit{A}}}\vDash\mathrm{(s\text{-}rank_{n})}$ or ${\boldsymbol{\mathit{A}}}\vDash\mathrm{(div_{p,n})}$.
Proof.
Denote
[…]
Since $\boldsymbol{\mathit{A}}$} is totally ordered we know that for each evaluation $e$ of $\boldsymbol{\mathit{A}}$} we have ${\boldsymbol{\mathit{A}}}\vDash_{e}\mathrm{(mix_{p,n})}$ iff ${\boldsymbol{\mathit{A}}}\vDash_{e}\mathrm{(div_{p,n})}$ or ${\boldsymbol{\mathit{A}}}\vDash_{e}\mathrm{(\text{s-rank}_{n}(y/x))}$. Therefore if ${\boldsymbol{\mathit{A}}}\vDash\mathrm{(div_{p,n})}$ or ${\boldsymbol{\mathit{A}}}\vDash\mathrm{(\text{s-rank}_{n}(y/x))}$ then clearly ${\boldsymbol{\mathit{A}}}\vDash\mathrm{(mix_{p,n})}$. For the other implication assume ${\boldsymbol{\mathit{A}}}\nvDash\mathrm{(div_{p,n})}$ and ${\boldsymbol{\mathit{A}}}\nvDash\mathrm{(\text{s-rank}_{n}(y/x))}$. Then there are evaluations $e_{1},e_{2}$ such that ${\boldsymbol{\mathit{A}}}\nvDash_{e_{1}}\mathrm{(div_{p,n})}$ and ${\boldsymbol{\mathit{A}}}\nvDash_{e_{2}}\mathrm{(\text{s-rank}_{n}(y/x))}$. We will define an evaluation $e_{3}$ on $\boldsymbol{\mathit{A}}$}, such that $e_{3}(x)=e_{1}(x)$ and $e_{3}(y)=e_{2}(y)$. Then we get ${\boldsymbol{\mathit{A}}}\nvDash_{e_{3}}\mathrm{(mix_{p,n})}$. Therefore we have ${\boldsymbol{\mathit{A}}}\vDash\mathrm{(mix_{p,n})}$ iff ${\boldsymbol{\mathit{A}}}\vDash\mathrm{(div_{p,n})}$ or ${\boldsymbol{\mathit{A}}}\vDash\mathrm{(\text{s-rank}_{n}(y/x))}$. Since ${\boldsymbol{\mathit{A}}}\vDash\mathrm{(\text{s-rank}_{n}(y/x))}$ if and only if ${\boldsymbol{\mathit{A}}}\vDash\mathrm{(\text{s-rank}_{n})}$, we obtain that ${\boldsymbol{\mathit{A}}}\vDash\mathrm{(mix_{p,n})}$ if and only if ${\boldsymbol{\mathit{A}}}\vDash\mathrm{(div_{p,n})}$ or ${\boldsymbol{\mathit{A}}}\vDash\mathrm{(\text{s-rank}_{n})}$, which completes the proof. ∎
Corollary 5.6.
Let $0\leq p\leq n$ and $m\leq n$. We have ${\boldsymbol{\mathit{Z}}}_{m}\vDash\mathrm{(mix_{p,n})}$. Moreover, ${\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(mix_{p,n})}$ iff $p\notin\operatorname{div}(m)$ or $m\leq 0$.
At this point we have all we need to provide the axiomatization of subvarieties of $p\mathbb{AL}$. First we will discuss subvarieties of positively and negatively pointed Abelian $\ell$-groups.
Theorem 5.7.
Any proper subvariety of $p\mathbb{AL}^{+}$ is of the form
$V_{I,J}=\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{i},{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\mid i\in I,j\in J)$
for some finite sets $J\subseteq I\subsetneq\mathbb{N}$, where both $I$ and $J$ are closed under divisors.
Moreover, $V_{I,J}$ is axiomatized by the following set of equations:
$S_{I,J}=\{\mathrm{(rank_{n})}\}\cup\{\mathrm{(div_{p,n})}\mid p\notin I\}\cup\{\mathrm{(mix_{p,n})}\mid p\in I\setminus J\},$
where $n=\max I$.
Proof.
First we show that any proper subvariety of $p\mathbb{AL}^{+}$ is equal to $V_{I,J}$ for some finite sets $J\subseteq I$.
Let $\mathbb{K}$}} be an arbitrary subvariety of $p\mathbb{AL}^{+}$. Let us denote $I=\{i\mid{\boldsymbol{\mathit{Z}}}_{i}\in{{\mathbb{K}}}\}$ and $J=\{j\mid{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\in{{\mathbb{K}}}\}$. We show ${{\mathbb{K}}}=V_{I,J}$. Since $\mathbb{K}$}} is a proper subvariety of $p\mathbb{AL}^{+}$, by Lemma 3.7 we know that ${\boldsymbol{\mathit{R}}}_{1}\notin{{\mathbb{K}}}$. By Theorem 4.10 we have
[…]
Clearly, both $I$ and $J$ are closed under divisors, since for any $d,k\in\mathbb{N}\setminus\{0\}$ such that $d\in\operatorname{div}(k)$ we have ${\boldsymbol{\mathit{Z}}}_{d}\in\mathbf{IS}({\boldsymbol{\mathit{Z}}}_{k})$ and ${\boldsymbol{\mathit{Z}}}_{d}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\in\mathbf{IS}({\boldsymbol{\mathit{Z}}}_{k}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$. Since for each $j\in J$ we have ${\boldsymbol{\mathit{Z}}}_{j}\in\mathbf{H}({\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$, we get $J\subseteq I$. The set $I$ (and consequently $J$) has to be finite, otherwise we would get by Lemma 4.3 that ${\boldsymbol{\mathit{R}}}_{1}\in{{\mathbb{K}}}$. This proves ${{\mathbb{K}}}=V_{I,J}$ for some finite sets $J\subseteq I$.
Now we have to show the axiomatization of $V_{I,J}$. First, we will argue that all the equations from $S_{I,J}$ hold in $V_{I,J}$. Since $n=\max I\geq\max J$, by Lemma 5.4 we know that ${\boldsymbol{\mathit{Z}}}_{i}\vDash\mathrm{(rank_{n})}$ for all $i\in I$ and ${\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(rank_{n})}$ for all $j\in J$. Thus $V_{I,J}\vDash\mathrm{(rank_{n})}$.
Let us fix $p\notin I$. Since $I$ is closed under divisors, we get that $i\notin\operatorname{div}(p)$ for all $i\in I$. Using $J\subseteq I$ and by Lemma 5.2 we obtain that ${\boldsymbol{\mathit{Z}}}_{i}\vDash\mathrm{(div_{p,n})}$ for each $i\in I$ and ${\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(div_{p,n})}$ for each $j\in J$. Therefore, $V_{I,J}\vDash\mathrm{(div_{p,n})}$ for each $p\notin I$.
Now let us fix $p\in I\setminus J$. Since $p\notin J$ by the same argument as in the previous paragraph we obtain ${\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(div_{p,n})}$ for each $j\in J$. By Lemma 5.1 we have ${\boldsymbol{\mathit{Z}}}_{i}\vDash\mathrm{(\text{s-rank}_{n})}$ for each $i\in I$ and thus by Lemma 5.5 we get ${\boldsymbol{\mathit{Z}}}_{i}\vDash\mathrm{(mix_{p,n})}$ and ${\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(mix_{p,n})}$ for each $i\in I$ and $j\in J$. Thus $V_{I,J}\vDash\mathrm{(mix_{p,n})}$ for each $p\in I\setminus J$.
This concludes that $V_{I,J}\vDash\mathrm{(eq)}$ for each $\mathrm{(eq)}\in S_{I,J}$. It remains to show $V_{I,J}$ is uniquely determined by the equations from $S_{I,J}$.
Let us have a finitely generated ${\boldsymbol{\mathit{A}}}_{a}\in p\mathbb{AL}^{+}_{SI}$ satisfying all the equations from $S_{I,J}$. We will show ${\boldsymbol{\mathit{A}}}_{a}\in V_{I,J}$. First, let us note that clearly $\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a})\neq\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{1})$ since from Lemma 3.7 we know $\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{1})=p\mathbb{AL}^{+}$ and $\boldsymbol{\mathit{A}}$} satisfies some equations, which are non-trivial in $p\mathbb{AL}^{+}$. By Theorems 3.10 and 4.10 it’s enough to consider cases ${\boldsymbol{\mathit{A}}}_{a}\in\{{\boldsymbol{\mathit{Z}}}_{i},{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\mid i,j\in\mathbb{N}\}$.
For ${\boldsymbol{\mathit{A}}}_{a}={\boldsymbol{\mathit{Z}}}_{m}$ we get $i\leq n$, by applying Lemma 5.1 using ${\boldsymbol{\mathit{Z}}}_{m}\vDash\mathrm{(rank_{n})}$. Since ${\boldsymbol{\mathit{Z}}}_{m}\vDash\mathrm{(div_{p,n})}$ for all $p\notin I$ we get by Lemma 5.2 that $m\in I$.
For ${\boldsymbol{\mathit{A}}}_{a}={\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}$ we similarly get $m\leq n$ and $m\in I$. Since for every $p\in I\setminus J$ we have ${\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(mix_{p,n})}$, by Lemma 5.5 we get $m\in J$.
Thus ${\boldsymbol{\mathit{A}}}_{a}\in V_{I,J}$, which completes the proof. ∎
The structural part of Theorem 5.7 was proved in [28]. This result can be obtained by applying the Mundici functor, Theorem 3.10, and Komori classification, an approach we will discuss later in Section 6 and demonstrate in Theorems 6.3 and 6.6.
Despite the fact that the statement of Theorem 5.7 was about proper varieties, we can also claim that any variety of $p\mathbb{AL}^{+}$ is equal to $V_{I,J}$ for some sets $J\subseteq I$, since $p\mathbb{AL}^{+}=V_{\mathbb{N},\mathbb{N}}$.
In this chapter, we have so far been describing the subvarieties of $p\mathbb{AL}^{+}$. However, we could state all the results similarly for $p\mathbb{AL}^{-}$ with accordingly modified equations:
[…]
Using these equations we can state a dual result to Theorem 5.7.
Theorem 5.8.
Any proper subvariety of $p\mathbb{AL}^{-}$ is of the form
$V^{d}_{I,J}=\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{i},{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\mid i\in I,j\in J)$
for some finite sets $J\subseteq I$.
Moreover, $V^{d}_{I,J}$ is axiomatized by the following set $S_{I,J}^{d}$ of equations:
$S_{I,J}^{d}=\{\mathrm{(rank_{n}^{d})}\}\cup\{\mathrm{(div_{p,n}^{d})}\mid p\notin I\}\cup\{\mathrm{(mix_{p,n}^{d})}\mid p\in I\setminus J\},$
where $n=-\min I$.
Now we can combine the Theorems 5.8 and 5.7 into one theorem covering all subvarieties of $p\mathbb{AL}$.
Theorem 5.9.
Any subvariety of $p\mathbb{AL}$ can be written as ${{\mathbb{K}}}=V_{I_{1},J_{1}}\lor V^{d}_{I_{2},J_{2}}$ for some (not necessarily finite) sets $I_{1},I_{2},J_{1},J_{2}$. Such a variety is axiomatized by the equations from $S_{I_{1},J_{1}}\cup S_{I_{2},J_{2}}^{d}$, where $S_{I_{1},J_{1}}$ and $S^{d}_{I_{2},J_{2}}$ are defined in Theorems 5.7 and 5.8.
Proof.
Let $\mathbb{K}$}} be arbitrary subvariety of $p\mathbb{AL}$. Using Theorem 2.5 we can derive that subdirectly irreducible pointed Abelian $\ell$-groups are totally ordered. Let us denote ${{\mathbb{K}}}_{SI}^{+}={{\mathbb{K}}}_{SI}\cap p\mathbb{AL}^{+}$ and ${{\mathbb{K}}}_{SI}^{-}={{\mathbb{K}}}_{SI}\cap p\mathbb{AL}^{-}$. Thus we have
$\mathbf{HSP}({{\mathbb{K}}}_{SI}^{+}\cup{{\mathbb{K}}}_{SI}^{-})=\mathbf{HSP}({{\mathbb{K}}}_{SI})={{\mathbb{K}}}.$
This shows that $\mathbb{K}$}} is the join of $\mathbf{HSP}({{\mathbb{K}}}_{SI}^{+})$ and $\mathbf{HSP}({{\mathbb{K}}}_{SI}^{-})$. By Theorem 5.7 $\mathbf{HSP}({{\mathbb{K}}}_{SI}^{+})=V_{I_{1},J_{1}}$ for some sets $I_{1},J_{1}$ and by Theorem 5.8 $\mathbf{HSP}({{\mathbb{K}}}_{SI}^{-})=V^{d}_{I_{2},J_{2}}$ for some sets $I_{2},J_{2}$. Thus we have ${{\mathbb{K}}}=V_{I_{1},J_{1}}\lor V^{d}_{I_{2},J_{2}}$.
We show that the equations from $S_{I_{1},J_{1}}\cup S_{I_{2},J_{2}}^{d}$ axiomatize the variety $\mathbb{K}$}}. First all the equations from $S_{I_{1},J_{1}}$ are valid in all negative pointed Abelian $\ell$-group by Lemmas 5.2, 5.4 and 5.5. Similarly, all the equations from $S^{d}_{I_{2},J_{2}}$ are valid in all positively pointed Abelian $\ell$-groups. Since all the equations from $S_{I_{1},J_{1}}$ are valid in $V_{I_{1},J_{1}}$ and all the equations from $S^{d}_{I_{2},J_{2}}$ are valid in $V^{d}_{I_{1},J_{1}}$ we conclude that all equations from $S_{I_{1},J_{1}}\cup S_{I_{2},J_{2}}^{d}$ are valid in $\mathbb{K}$}}.
Now let us take arbitrary ${\boldsymbol{\mathit{A}}}_{a}\in p\mathbb{AL}_{SI}$. By Theorem 2.5 ${\boldsymbol{\mathit{A}}}_{a}$ is totally ordered. We distinguish three cases:
  1. 1.
    If $a=0$ then ${\boldsymbol{\mathit{A}}}_{a}\in{{\mathbb{K}}}$ by Corollary 3.6.
  2. 2.
    If $a>0$ then ${\boldsymbol{\mathit{A}}}_{a}\in{{\mathbb{K}}}_{SI}^{+}$ and thus ${\boldsymbol{\mathit{A}}}\in V_{I_{1},J_{1}}$. Consequently ${\boldsymbol{\mathit{A}}}\in{{\mathbb{K}}}$.
  3. 3.
    If $a<0$ then ${\boldsymbol{\mathit{A}}}_{a}\in{{\mathbb{K}}}_{SI}^{-}$ and thus ${\boldsymbol{\mathit{A}}}\in V^{d}_{I_{2},J_{2}}$. This implies ${\boldsymbol{\mathit{A}}}\in{{\mathbb{K}}}$.
This proves that the equations from $S_{I_{1},J_{1}}\cup S_{I_{2},J_{2}}^{d}$ axiomatize $\mathbb{K}$}}. ∎
6 Applying Mundici functor
In the previous sections, we developed a self-contained theory using the language of Abelian $\ell$-groups. In this section we will establish several connections between theory of MV-algebras and theory of pointed Abelian $\ell$-groups. We will show that the semantical description of subvarieties of $\mathrm{pAb}$$^{+}$ is mutually equivalent with the Komori classification of subvarieties of MV-algebras. We will conclude the section by deriving a complete classification of quasivarieties of $\mathrm{pAb}$} generated by totally ordered elements from similar result about MV-algebras [14].
Let us recall that MV-algebras can be understood as bounded intervals of Abelian $\ell$-groups. For us, MV-algebra will be a structure of the following signature $\langle{\oplus,\otimes,\lor,\land,\neg,0,1}\rangle$. For the basics of MV-algebra we recommend the reader to check any classical bibliography e.g. [7].
We shall note that we are adding the symbols $\lor,\land$ and $\otimes$ into our language purely for our convenience, it is well-known that $\lor,\land$ and $\otimes$ are term definable using the remaining operations.
The main tool we use in this section is Mundici functor, which is a functor from the category of strongly positively pointed Abelian $\ell$-groups into the category of MV-algebras. We will denote the Mundici functor by $\Gamma$. Here, we provide the definition of $\Gamma$ on objects.
Definition 6.1.
Let ${\boldsymbol{\mathit{A}}}_{u}=\langle{A,+,-,\lor,\land,0,u}\rangle$ be a positively pointed Abelian $\ell$-group and $u\geq 0$. Define MV-algebra $\Gamma({\boldsymbol{\mathit{A}}}_{u})=\langle{[0,u],\oplus,\odot,\lor,\land,\neg,0,u}\rangle$, where the operations are defined as follows:
$a\oplus b=(a+b)\land u$
$a\otimes b=0\lor(a+b-u)$
$\neg a=u-a$
Similarly, for a homomorphism $f:{\boldsymbol{\mathit{A}}}_{u}\rightarrow{\boldsymbol{\mathit{B}}}_{v}$ we define $\Gamma(f):\Gamma({\boldsymbol{\mathit{A}}}_{u})\rightarrow\Gamma({\boldsymbol{\mathit{B}}}_{v})$ as $\Gamma(f)=f\restriction[0,u]$. It is well-known that $\Gamma$ is indeed a functor. Moreover, it is known [7] that $\Gamma$ is a categorical equivalence. We will denote its ”inverse” by $\Gamma^{-1}$.
It can be elementarily shown (as an exercise or see [28]) that $\Gamma$ and $\Gamma^{-1}$ preserve products, subalgebras and homomorphisms. To be more precise, we state this lemma.
Lemma 6.2.
For any positively strongly pointed Abelian $\ell$-groups ${\boldsymbol{\mathit{A}}}_{u},{\boldsymbol{\mathit{B}}}_{v}$ the following hold:
  1. 1.
    $\Gamma(\prod_{i\in I}{\boldsymbol{\mathit{A}}}_{i})=\prod_{i\in I}\Gamma({\boldsymbol{\mathit{A}}}_{i})$,
  2. 2.
    ${\boldsymbol{\mathit{A}}}_{u}\in\mathbf{IS}({\boldsymbol{\mathit{B}}}_{v})$ iff $\Gamma({\boldsymbol{\mathit{A}}}_{u})\in\mathbf{IS}(\Gamma({\boldsymbol{\mathit{B}}}_{v}))$,
  3. 3.
    ${\boldsymbol{\mathit{A}}}_{u}\in\mathbf{H}({\boldsymbol{\mathit{B}}}_{v})$ iff $\Gamma({\boldsymbol{\mathit{A}}}_{u})\in\mathbf{H}(\Gamma({\boldsymbol{\mathit{B}}}_{v}))$.
Using this observation we can easily provide the syntactical classification of all varieties of MV-algebras.
Theorem 6.3 (Komori).
Every variety of MV-algebras is generated by
$\{\Gamma({\boldsymbol{\mathit{Z}}}_{i}),\Gamma({\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})\mid i\in I,j\in J\}\text{ for some finite sets }J\subseteq I.$
Proof.
Let $\mathbb{K}$}} be a proper variety of MV-algebras. Denote some ${\boldsymbol{\mathit{A}}}\notin{{\mathbb{K}}}$. Then we have by the lemma above $\Gamma^{-1}({\boldsymbol{\mathit{A}}})\notin\mathbf{HSP}(\Gamma^{-1}[{{\mathbb{K}}}])$. Thus $\mathbf{HSP}(\Gamma^{-1}[{{\mathbb{K}}}])$ is a proper subvariety of $p\mathbb{AL}^{+}$ and thus we have $\mathbf{HSP}(\Gamma^{-1}[{{\mathbb{K}}}])=V_{I,J}$ for some finite sets $J\subseteq I$. Since ${\boldsymbol{\mathit{Z}}}_{i},{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\in V_{I,J}$ for $i\in I,j\in J$, by lemma above we obtain $\Gamma({\boldsymbol{\mathit{Z}}}_{i}),\Gamma({\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})\in{{\mathbb{K}}}$ for $i\in I$ and $j\in J$. Thus $\{\Gamma({\boldsymbol{\mathit{Z}}}_{i}),\Gamma({\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})\mid i\in I,j\in J\}\subseteq{{\mathbb{K}}}.$
To show the other inclusion we assume that ${\boldsymbol{\mathit{A}}}\in{{\mathbb{K}}}$. Consequently, $\Gamma^{-1}({\boldsymbol{\mathit{A}}})\in\Gamma^{-1}[{{\mathbb{K}}}]$ and thus $\Gamma^{-1}({\boldsymbol{\mathit{A}}})\in V_{I,J}$ which means ${\boldsymbol{\mathit{A}}}\in\{\Gamma({\boldsymbol{\mathit{Z}}}_{i}),\Gamma({\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})\mid i\in I,j\in J\}$. This proves ${{\mathbb{K}}}=\{\Gamma({\boldsymbol{\mathit{Z}}}_{i}),\Gamma({\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})\mid i\in I,j\in J\}$ for some finite sets $J\subseteq I$. ∎
Similarly, one can show semantical part of Theorem 5.9 using Komori classification and Theorem 3.10. For details one can see discussion in [28]. We prove here something stronger.
Let $\mathbf{S}$ denote any finitely generated dense $\ell$-subgroup of $\mathbf{R}$ such that $\mathbf{S}\cap\mathbf{Q}=\mathbf{Z}$. Recall the result from [14]:
Theorem 6.4.
Let $\mathbb{M}$ be a class of MV-chains. Then the following properties are equivalent:
  1. 1.
    $\mathbb{M}$ is universal.
  2. 2.
    There exist $I,J,K\subseteq\mathbb{N}$ and for every $j\in J$, a nonempty subset $D_{j}\subseteq\operatorname{div}(j)$ such that $\mathbb{M}$ is equal to
    $\mathbf{ISP}_{\mathrm{U}}(\{\Gamma({\boldsymbol{\mathit{Z}}}_{i})\mid i\in I\}\cup\{\Gamma({\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{d_{j}})\mid j\in J,d_{j}\in D_{j}\}\cup\{\Gamma(\mathbf{S}_{k})\mid k\in K\}).$
We need to show how Mundici functor behaves with respect to ultrapowers. Generally, it is not true that $\Gamma(\prod_{i\in\omega}{\boldsymbol{\mathit{A}}}_{a}/\mathfrak{U})=\prod_{i\in\omega}\Gamma({\boldsymbol{\mathit{A}}}_{a})/\mathfrak{U}$, where $\mathfrak{U}$ is a non-principal ultrafilter. In fact one can easily show that for strongly pointed $\ell$-group ${\boldsymbol{\mathit{A}}}_{a}$ the $\ell$-group $\prod_{i\in\omega}{\boldsymbol{\mathit{A}}}_{a}/\mathfrak{U}$ is never even strongly pointed. However, we claim that Mundici functor preserves universal subclasses. We will need the following lemma, which we prove later in the end of this section.
Lemma 6.5.
Let ${{\mathbb{K}}}\cup\{{\boldsymbol{\mathit{A}}}_{u}\}$ be a class of totally ordered strongly positively pointed Abelian $\ell$-groups. Then we have ${\boldsymbol{\mathit{A}}}_{u}\in\mathbf{ISP}_{\mathrm{U}}({{\mathbb{K}}})$ iff $\Gamma({\boldsymbol{\mathit{A}}}_{u})\in\mathbf{ISP}_{\mathrm{U}}(\Gamma[{{\mathbb{K}}}])$.
Using these tools we can prove the following.
Theorem 6.6.
Let $\mathbb{M}$ be a class of totally ordered positively pointed Abelian $\ell$-groups. Then the following properties are equivalent:
  1. 1.
    $\mathbb{M}$ is universal.
  2. 2.
    There exist $I,J,K\subseteq\mathbb{N}$ and for every $j\in J$, a nonempty subset $D_{j}\subseteq\operatorname{div}(j)$ such that
    $\mathbb{M}=\mathbf{ISP}_{\mathrm{U}}(\{{\boldsymbol{\mathit{Z}}}_{i}\mid i\in I\}\cup\{{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{d_{j}}\mid j\in J,d_{j}\in D_{j}\}\cup\{\mathbf{S}_{k}\mid k\in K\}).$
Proof.
Implication $2$ to 1 is trivial. To prove the other implication assume $\mathbb{M}$ is universal. Therefore, by Lemma 2.3 we have $\mathbb{M}=\mathbf{ISP}_{\mathrm{U}}(\mathbb{M})$. Let us denote $\mathbb{K}$}} the subclass of $M$ consisting of all strongly pointed Abelian $\ell$-groups.
Using Lemma 6.5 it follows $\Gamma[{{\mathbb{K}}}]=\mathbf{ISP}_{\mathrm{U}}(\Gamma[{{\mathbb{K}}}])$. Thus $\Gamma[{{\mathbb{K}}}]$ is universal and hence by Theorem 6.4 we have that there exist $I,J,K\subseteq\mathbb{N}$ and for every $j\in J$, a nonempty subset $D_{j}\subseteq\operatorname{div}(j)$ such that
$\Gamma[{{\mathbb{K}}}]=\mathbf{ISP}_{\mathrm{U}}(\{\Gamma({\boldsymbol{\mathit{Z}}}_{i})\mid i\in I\}\cup\{\Gamma({\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{d_{j}})\mid j\in J,d_{j}\in D_{j}\}\cup\{\Gamma(\mathbf{S}_{k})\mid k\in K\}).$
Now, again using Lemma 6.5 we obtain
${{\mathbb{K}}}=\mathbf{ISP}_{\mathrm{U}}(\{{\boldsymbol{\mathit{Z}}}_{i}\mid i\in I\}\cup\{{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{d_{j}}\mid j\in J,d_{j}\in D_{j}\}\cup\{\mathbf{S}_{k}\mid k\in K\})\cap sp\mathbb{AL}^{+}.$
We discuss two options:
  1. 1.
    If ${\boldsymbol{\mathit{Z}}}_{0}\in\mathbb{M}$ then $p\mathbb{AL}_{SI}^{0}\subseteq\mathbb{M}$ by Lemma 2.11 and thus by Theorem 3.10 we have
    $\mathbb{M}=\mathbf{ISP}_{\mathrm{U}}({{\mathbb{K}}})\cup p\mathbb{AL}_{SI}^{0}=\mathbf{ISP}_{\mathrm{U}}({{\mathbb{K}}})\cup\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{0})=$
    $\mathbf{ISP}_{\mathrm{U}}(\{{\boldsymbol{\mathit{Z}}}_{i}\mid i\in I\}\cup\{{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{d_{j}}\mid j\in J,d_{j}\in D_{j}\}\cup\{\mathbf{S}_{k}\mid k\in K\}).$
  2. 2.
    If $Z_{0}\notin\mathbb{M}$, we get by Lemma 2.11 that $p\mathbb{AL}_{SI}^{0}\cap\mathbb{M}=\emptyset$ and thus $\mathbb{M}=$
$\mathbf{ISP}_{\mathrm{U}}({{\mathbb{K}}})=\mathbf{ISP}_{\mathrm{U}}(\{{\boldsymbol{\mathit{Z}}}_{i}\mid i\in I\}\cup\{{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{d_{j}}\mid j\in J,d_{j}\in D_{j}\}\cup\{\mathbf{S}_{k}\mid k\in K\}).$
This completes the proof. ∎
Corollary 6.7.
Let $\mathbb{K}$}} be the quasivariety generated by a class of totally ordered positively pointed Abelian $\ell$-groups. Then there exists $I,J,K\subseteq\mathbb{N}$ and for every $j\in J$, a nonempty subset $D_{j}\subseteq\operatorname{div}(j)$ such that
${{\mathbb{K}}}=\mathbf{ISPP}_{\mathrm{U}}(\{{\boldsymbol{\mathit{Z}}}_{i}\mid i\in I\}\cup\{{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{d_{j}}\mid j\in J,d_{j}\in D_{j}\}\cup\{\mathbf{S}_{k}\mid k\in K\}).$
Let us just say that Theorem 6.6 and Corollary 6.7 can be stated also for negatively pointed Abelian $\ell$-groups and thus we can get description of all universal classes of totally ordered pointed Abelian $\ell$-groups. This gives us two corollaries.
Corollary 6.8.
Let $\mathbb{M}$ be a class of totally ordered positively pointed Abelian $\ell$-groups. Then the following properties are equivalent:
  1. 1.
    $\mathbb{M}$ is universal.
  2. 2.
    There exist $I,J,K\subseteq\mathbb{Z}$ and for every $j\in J$, a nonempty subset $D_{j}\subseteq\operatorname{div}(j)$ such that
    $\mathbb{M}=\mathbf{ISP}_{\mathrm{U}}(\{{\boldsymbol{\mathit{Z}}}_{i}\mid i\in I\}\cup\{{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{d_{j}}\mid j\in J,d_{j}\in D_{j}\}\cup\{\mathbf{S}_{k}\mid k\in K\}).$
Corollary 6.9.
Let $\mathbb{K}$}} be the quasivariety generated by a class of totally ordered positively pointed Abelian $\ell$-groups. Then there exists $I,J,K\subseteq\mathbb{N}$ and for every $j\in J$, a nonempty subset $D_{j}\subseteq\operatorname{div}(j)$ such that
${{\mathbb{K}}}=\mathbf{ISPP}_{\mathrm{U}}(\{{\boldsymbol{\mathit{Z}}}_{i}\mid i\in I\}\cup\{{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{d_{j}}\mid j\in J,d_{j}\in D_{j}\}\cup\{\mathbf{S}_{k}\mid k\in K\})$
In the rest of the section we will focus on proving Lemma 6.5. Before doing so, we need a few smaller lemmas.
Lemma 6.10.
Let ${\boldsymbol{\mathit{A}}}_{u}$ be a totally ordered strongly positively pointed Abelian $\ell$-group. Then for each $a\in A$ there is a unique $n\in\mathbb{Z}$ and unique $r\in A$ such that $0\leq r<u$ and $a=n\cdot u+r$.
Proof.
Since $u$ is strong unit of $\boldsymbol{\mathit{A}}$} there exist $n,m\in\mathbb{Z}$ such that $n\cdot u\leq a<m\cdot u$. Let us fix maximum such $n$. Then we have $n\cdot u\leq a<(n+1)\cdot u$ and $0\leq a-n\cdot u<u$. Let us set $r=a-n\cdot u$. We indeed have $n\cdot u+r=a$.
We prove uniqueness of such decomposition. Assume $a=n_{1}\cdot u+r_{1}=n_{2}\cdot u+r_{2}$ for some $n_{1},n_{2}\in\mathbb{Z}$ and $r_{1},r_{2}\in A$ such that $0\leq r_{1},r_{2}<u$. We get $(n_{1}-n_{2})\cdot u=r_{2}-r_{1}$. Since $u<r_{2}-r_{1}<u$ we obtain $n_{1}=n_{2}$ and $r_{1}=r_{2}$. This completes the proof. ∎
The following observation directly follows from the definition of $\oplus$ and $\otimes$.
Lemma 6.11.
Let ${\boldsymbol{\mathit{A}}}_{u}$ be a positively pointed Abelian $\ell$-group and $0\leq a,b\leq u$.
  1. 1.
    If $a+b\leq u$ or $a\oplus b<u$ then $a+b=a\oplus b$.
  2. 2.
    If $a+b\geq u$ or $a\otimes b>0$ then $a+b-u=a\otimes b$.
Lemma 6.12.
Let $\boldsymbol{\mathit{A}}$} be a totally ordered pointed Abelian $\ell$-group and let $F$ be a finite subset of $A$. Let $f_{F}$ be a partial embedding of $F$ into a class of pointed Abelian $\ell$-groups $\mathbb{K}$}}. Let $a,b\in F$ such that $a\leq b$. Then $f_{F}(a)\leq f_{F}(b)$. Moreover, if $a\neq b$, then $f_{F}(a)<f_{F}(b)$.
Proof.
Since $a\lor b=b$ and $a,b,a\lor b\in F$ we have by definition of partial embedding that $f_{F}(a)\lor f_{F}(b)=f_{F}(a\lor b)$, thus $f_{F}(a)\leq f_{F}(a\lor b)=f_{F}(b)$. Also, whenever $a\neq b$ then, since $f_{F}$ is injective, we obtain $f_{F}(a)<f_{F}(b)$. ∎
Now, we can finally prove the Lemma 6.5.
Proof of Lemma 6.5.
Using Lemma 2.3 we have to show that ${\boldsymbol{\mathit{A}}}_{u}$ is partially embeddable into $\mathbb{K}$}} iff $\Gamma({\boldsymbol{\mathit{A}}}_{u})$ is partially embeddable into $\Gamma[{{\mathbb{K}}}]$.
First assume that for strongly pointed $\ell$-group ${\boldsymbol{\mathit{A}}}_{u}$ there is a family of partial embeddings $\{f_{F}\}_{F\subseteq A,|F|<\omega}$, where $f_{F}$ is a partial embedding of the set $F$ into some ${\boldsymbol{\mathit{B}}}_{v}\in{{\mathbb{K}}}$.
Let $G\subseteq\Gamma({\boldsymbol{\mathit{A}}}_{u})$ be an arbitrary finite set. Let us define $H=\{g_{1}+g_{2},-g_{1}\mid g_{1},g_{2}\in G\}$. We show that the mapping $f_{H}$ is a partial embedding of $G$ into $\Gamma({\boldsymbol{\mathit{B}}}_{v})\in\Gamma[{{\mathbb{K}}}]$. We already know that $f_{H}$ is injective and it preserves constants. Since $\otimes,\lor$ and $\land$ are definable in MV-algebras using the other connectives it is enough to show that $f_{H}$ preserves $\oplus$ and $\neg$.
We show $f_{H}$ preserves $\oplus$. Let us assume $a_{1},a_{2},a_{1}\oplus a_{2}\in G$. Let us consider two cases:
  1. 1.
    Assume $a_{1}+a_{2}\geq u$. Then we have by Lemma 6.12 $f_{H}(a_{1}+a_{2})\geq f_{H}(u)=v$ and thus
    $f_{H}(a_{1}\oplus a_{2})=f_{H}((a_{1}+a_{2})\land u)=f_{H}(u)=v=\\ f_{H}(a_{1}+a_{2})\land v=f_{H}(a_{1})+f_{H}(a_{2})\land v=f_{H}(a_{1})\oplus f_{H}(a_{2}).$
  2. 2.
    Now assume $a_{1}+a_{2}\leq u$. Then we have by Lemma 6.12 $f_{H}(a_{1}+a_{2})\leq f_{H}(u)=v$
    We have
    $f_{H}(a_{1}\oplus a_{2})=f_{H}((a_{1}+a_{2})\land u)=f_{H}(a_{1}+a_{2})=\\ f_{H}(a_{1})+f_{H}(a_{2})=(f_{H}(a_{1})+f_{H}(a_{2}))\land v=f_{H}(a_{1})\oplus f_{H}(a_{2}).$
It remains to check $f_{H}$ preserves $\neg$. Let us assume $a,\neg a\in G$. Then $a\in H$ and we have
$f_{H}(\neg a)=f_{H}(u-a)=f_{H}(u)+f_{H}(-a)=v-f_{H}(a)=\neg f_{H}(a).$
Therefore, $f_{H}$ preserves $\neg$. This shows $f_{H}$ is indeed a partial embedding of $G$ from $\Gamma({\boldsymbol{\mathit{A}}})$ into $\Gamma({\boldsymbol{\mathit{B}}})$. Since $G$ was arbitrary set we obtain that $\Gamma({\boldsymbol{\mathit{A}}})$ is partially embeddable into $\Gamma({\boldsymbol{\mathit{B}}})$.
Now, we will prove that if $\Gamma({\boldsymbol{\mathit{A}}}_{u})$ is partially embeddable into $\Gamma({\boldsymbol{\mathit{B}}}_{v})$ then also ${\boldsymbol{\mathit{A}}}_{u}$ is partially embeddable into ${\boldsymbol{\mathit{B}}}_{v}$.
Assume that $\Gamma({\boldsymbol{\mathit{A}}}_{u})$ is partially embeddable into $\Gamma({\boldsymbol{\mathit{B}}}_{v})$. Therefore for any finite set $S\subseteq\Gamma({\boldsymbol{\mathit{A}}}_{u})$ there is partial embedding $g_{S}:S\rightarrow\Gamma({\boldsymbol{\mathit{B}}}_{v})$.
Let $F\subseteq A_{u}$ be finite. Let us define a finite superset $F^{\prime}$, which is closed under - and also that $(n\cdot u+a)\in F^{\prime}\Rightarrow(n\cdot u)\in F^{\prime}\land a\in F^{\prime}$ for $n\in{\boldsymbol{\mathit{Z}}}$ and $a\in[0,u)$. We denote $F_{0}=F^{\prime}\cap[0,u]$. For any $a\in A$ we denote the decomposition from Lemma 6.10 $a=n_{a}\cdot u+r_{a}$, where $n_{a}\in Z$ and $r_{a}\in[0,u)$. Since $F_{0}\subseteq\Gamma({\boldsymbol{\mathit{A}}}_{u})$ we can define a partial a embedding $f_{F}\rightarrow B_{v}$ as follows:
$f_{F}(a)=(n_{a}\cdot v+g_{F_{0}}(r_{a})).$
This mapping is clearly injective, since we can construct $a$ knowing $n_{a},r_{a}$. We will verify it is indeed a partial homomorphism. Let $a,b,a+b\in F$. We have
$f_{F}(a+b)=n_{a+b}\cdot v+g_{F_{0}}(r_{a+b}).$
Now let us distinguish two cases.
  1. 1.
    First let us consider the case $r_{a}+r_{b}<u$ and thus by Lemma 6.11 we obtain $r_{a+b}=r_{a}+r_{b}=r_{a}\oplus r_{b}$ and $n_{a+b}=n_{a}+n_{b}$. By Lemma 6.12 and 6.11 we have
    $v>g_{F_{0}}(r_{a}\oplus r_{b})=g_{F_{0}}(r_{a})\oplus g_{F_{0}}(r_{b})=g_{F_{0}}(r_{a})+g_{F_{0}}(r_{b})$
    and thus
    $f_{F}(a+b)=n_{a}\cdot v+n_{b}\cdot v+g_{F_{0}}(r_{a}+r_{b})=n_{a}\cdot v+n_{b}\cdot v+g_{F_{0}}(r_{a}\oplus r_{b})=$
    $=n_{a}\cdot v+n_{b}\cdot v+g_{F_{0}}(r_{a})+g_{F_{0}}(r_{b})=$
    $n_{a}\cdot v+g_{F_{0}}(r_{a})+n_{b}\cdot v+g_{F_{0}}(r_{b})=f_{F}(a)+f_{F}(b).$
  2. 2.
    Now consider the case when $r_{a}+r_{b}>u$. Then we have $r_{a}+r_{b}=u+r_{a+b}$, $n_{a}+n_{b}+1=n_{a+b}$ and by Lemma 6.11 $r_{a}\otimes r_{b}=r_{a}+r_{b}-u$. By Lemma 6.12 and 6.11 we get
    $0<g_{F_{0}}(r_{a}\otimes r_{b})=g_{F_{0}}(r_{a})\otimes g_{F_{0}}(r_{b})=g_{F_{0}}(r_{a})+g_{F_{0}}(r_{b})-v$
    and thus
    $f_{F}(a+b)=(n_{a}+n_{b}+1)\cdot v+g_{F_{0}}(r_{a}+r_{b}-u)=$
    $(n_{a}+n_{b}+1)\cdot v+g_{F_{0}}(r_{a}\otimes r_{b})=$
    $(n_{a}+n_{b})\cdot v+v+g_{F_{0}}(r_{a})+g_{F_{0}}(r_{b})-v=$
    $n_{a}\cdot v+g_{F_{0}}(r_{a})+n_{b}\cdot v+g_{F_{0}}(r_{b})=f_{F}(a)+f_{F}(b).$
  3. 3.
    In the last case $r_{a}+r_{b}=u$ we have $r_{a+b}=0$, $n_{a+b}=n_{a}+n_{b}+1$ and $g_{F_{0}}(r_{a}+r_{b})=v$. Therefore,
    $f_{F}(a+b)=(n_{a}+n_{b}+1)\cdot v=$
    $n_{a}\cdot v+n_{b}\cdot v+g_{F_{0}}(r_{a}+r_{b})=n_{a}\cdot v+n_{b}\cdot v+g_{F_{0}}(r_{a}\oplus r_{b})=$
    $n_{a}\cdot v+g_{F_{0}}(r_{a})+n_{b}\cdot v+g_{F_{0}}(r_{b})=f_{F}(a)+f_{F}(b).$
This shows $f_{F}$ is preserving ++. Clearly, $f_{F}$ is preserving 0 and $a$. We will show $f_{F}$ is preserving -.
We have $f_{F}(a)=-(n_{a}\cdot v+g_{F_{0}}(r_{a}))=-n_{a}\cdot v-g_{F_{0}}(r_{a})$. If $r_{a}=0$ we have $f_{F}(a)=-n_{a}\cdot v=n_{-a}\cdot v=f_{F}(-a)$. Otherwise we have $n_{a}\cdot v-g_{F_{0}}(r_{a})<-n_{a}\cdot v$ and thus $n_{a}=n_{-a}+1$ and $u-r_{a}=r_{-a}$. By applying $g_{F_{0}}$ we get
$v-g_{F_{0}}(r_{a})=\neg g_{F_{0}}(r_{a})=g_{F_{0}}(\neg r_{a})=g_{F_{0}}(u-r_{a})=g_{F_{0}}(r_{-a}).$
Therefore, we have
$f_{F}(a)=-n_{a}\cdot v-g_{F_{0}}(r_{a})=n_{-a}\cdot v+v-v+g_{F_{0}}(r_{-a})=n_{-a}\cdot v+g_{F_{0}}(r_{-a})=f_{F}(-a).$
This shows $f_{F}$ is preserving -. It remains to check $f_{F}$ preserves lattice operations. Since $\boldsymbol{\mathit{A}}$} is totally ordered it is enough to check $f_{F}$ preserves ordering. Let $a\leq b$ and consequently $n_{a}\leq n_{b}$. We distinguish two cases
  1. 1.
    If $n_{a}=n_{b}$ then $r_{a}\leq r_{b}$. By Lemma 6.12, also $g_{F_{0}}(r_{a})\leq g_{F_{0}}(r_{b})$ and therefore we have the following:
    $f_{F}(a)=n_{a}\cdot v+g_{F_{0}}(r_{a})\leq n_{a}\cdot v+g_{F_{0}}(r_{b})=f_{F}(b).$
  2. 2.
    If $n_{a}<n_{b}$, by using Lemma 6.12 we obtain $g_{F_{0}}(r_{a})\leq g_{F_{0}}(u)=v$ and $0=g_{F_{0}}(0)\leq g_{F_{0}}(r_{b})$ and thus we have
    $f_{F}(a)=n_{a}\cdot v+g_{F_{0}}(r_{a})\leq(n_{a}+1)\cdot v\leq n_{b}\cdot v\leq n_{b}\cdot v+g_{F_{0}}(r_{b})=f_{F}(b).$
This shows $f_{F}$ preserves ordering and thus $f_{F}$ a partial embedding. Since the set $F$ was arbitrary, we conclude that ${\boldsymbol{\mathit{A}}}_{u}$ is partially embeddable into $\mathbb{K}$}}. ∎
$p\mathbb{AL}$$\mathcal{V}(\mathbf{Z}_{0})$$\mathcal{T}$$p\mathbb{AL}^{+}$$\mathcal{V}(\mathbf{Z}_{2})$$\mathcal{V}(\mathbf{Z}_{3})$$\mathcal{V}(\mathbf{Z}_{5})$$\mathcal{V}(\mathbf{Z}_{4})$$\mathcal{V}(\mathbf{Z}_{6})$$\mathcal{V}(\mathbf{Z}_{9})$$\mathcal{V}(\mathbf{Z}_{12})$$\mathcal{V}(\mathbf{Z}_{2}\overrightarrow{\times}\mathbf{Z}_{0})$$\mathcal{V}(\mathbf{Z}_{3}\overrightarrow{\times}\mathbf{Z}_{0})$$\mathcal{V}(\mathbf{Z}_{5}\overrightarrow{\times}\mathbf{Z}_{0})$$\mathcal{V}(\mathbf{Z}_{4}\overrightarrow{\times}\mathbf{Z}_{0})$$\mathcal{V}(\mathbf{Z}_{6}\overrightarrow{\times}\mathbf{Z}_{0})$$\mathcal{V}(\mathbf{Z}_{12}\overrightarrow{\times}\mathbf{Z}_{0})$$\mathcal{V}(\mathbf{Z}_{9}\overrightarrow{\times}\mathbf{Z}_{0})$$p\mathbb{AL}^{-}$$\mathcal{V}(\mathbf{Z}_{-2})$$\mathcal{V}(\mathbf{Z}_{-3})$$\mathcal{V}(\mathbf{Z}_{-5})$$\mathcal{V}(\mathbf{Z}_{-4})$$\mathcal{V}(\mathbf{Z}_{-6})$$\mathcal{V}(\mathbf{Z}_{-9})$$\mathcal{V}(\mathbf{Z}_{-12})$$\mathcal{V}(\mathbf{Z}_{-2}\overrightarrow{\times}\mathbf{Z}_{0})$$\mathcal{V}(\mathbf{Z}_{-3}\overrightarrow{\times}\mathbf{Z}_{0})$$\mathcal{V}(\mathbf{Z}_{-5}\overrightarrow{\times}\mathbf{Z}_{0})$$\mathcal{V}(\mathbf{Z}_{-4}\overrightarrow{\times}\mathbf{Z}_{0})$$\mathcal{V}(\mathbf{Z}_{-6}\overrightarrow{\times}\mathbf{Z}_{0})$$\mathcal{V}(\mathbf{Z}_{-12}\overrightarrow{\times}\mathbf{Z}_{0})$$\mathcal{V}(\mathbf{Z}_{-9}\overrightarrow{\times}\mathbf{Z}_{0})$
Figure 1: The lattice of subvarieties of $p\mathbb{AL}$.
[…]
Figure 2: The lattice of subvarieties of $p\mathbb{AL}$
Acknowledgements
This work was supported by the Czech Science Foundation grant 25-17958J. This work was also financially supported by Charles University Grant Agency, project no. 101724, entitled ”Exploring the foundations of reasoning for rational interaction under conditions of uncertainty”, implemented at the Faculty of Arts of Charles University.